### v1

# Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers

Abstract.
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine and human-authored code. Through a rigorous analysis of code attributes such as length, lexical diversity, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the structural segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct structural patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
copyright: none
1. Introduction
The advent of large language models (LLMs) such as Codex (Chen et al., 2021) and ChatGPT (OpenAI, 2022) has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect. LLMs trained on code (Allal et al., 2023; Nijkamp et al., 2023; Zheng et al., 2023; Rozière et al., 2023), sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation (Yetiştiren et al., 2023; Xu et al., 2022). Their ability to generate syntactically correct and functionally robust code challenges traditional dynamics and introduces a new era of efficiency and innovation in software creation, maintenance, and evolution.
While capable of generating human-like programs, LLMs bring ambiguity of whether a software artifact is created by a human or machine, causing integrity and authenticity issues in software development. This indistinction can lead to various challenges, such as the misattribution of code ownership for bug/fault triage and potential vulnerabilities in machine-generated code that may go unnoticed due to overreliance on its perceived robustness. Furthermore, the ease of generating code could lead to inflated assessments of project workloads, skewing economic valuations and potentially compromising the reliability of software. This blending of human and machine efforts not only raises questions about the trustworthiness of the software but also threatens the integrity of coding as a discipline, wherein the true authorship and the effort invested in creating software artifacts become obscured. Addressing these concerns is pivotal in maintaining a transparent and secure software development lifecycle.
Recently, there has been a growing research trend in detecting machine-generated texts (Yang et al., 2023; Zhao et al., 2023). Perturbation-based methods like DetectGPT (Mitchell et al., 2023), have achieved state-of-the-art results in identifying machine-generated text. DetectGPT employs likelihood score discrepancies between the original text and its perturbed variants to enhance detection accuracy. And the perturbation process is performed by a pretrained language model (e.g., T5 (Raffel et al., 2020)) to mask portions of the original text and then reconstruct it. While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid structural constraints. Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code (Lee et al., 2023).
In this paper, we conduct a comparative analysis of the patterns between machine and human-written code from five aspects, including length, lexical diversity, token frequency, token category proportions, and naturalness. Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces.
Capitalizing on our findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code. Our DetectCodeGPT improves DetectGPT by capturing the distinct structural patterns of code. Diverging from conventional techniques that depend on external pretrained model (e.g. CodeT5 (Wang et al., 2021)) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code corpus by strategically inserting spaces and newlines. Eschewing the need for an LLM for perturbations, our method provides a much higher efficiency, Since the proposed perturbation strategy well captures the structural distinctions between machine and human code, our DetectCodeGPT also ensures higher efficacy.
To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models (LLMs), ranging from 1.3 to 7 billion parameters, and operating at two different decoding temperatures. The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward (Yang et al., 2023; Tian et al., 2023), detecting machine-authored code presents a unique challenge. Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement with an average absolute increase of 6.46% in AUROC.
Our contributions can be summarized as follows:
  • To our knowledge, we are the first to conduct a comprehensive and thorough analysis of the differences between machine- and human-authored code. Our study sheds light on essential insights that further advance the utility of LLMs in coding.
  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct structural patterns of code.
  • We extensively evaluate the DetectCodeGPT across a variety of settings and show the effectiveness of our approach.
2. Background
2.1. Large Language Models for Code
GPT language models (Radford et al., 2018, 2019) based on Transformer (Vaswani et al., 2017) decoder has achieved remarkable success in natural language processing tasks (Ray, 2023). In the file of code generation, Codex (Chen et al., 2021) and AlphaCode (Li et al., 2022) are pioneering works to train large language models on code. These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens (Radford et al., 2018). Mathematically, this is captured as:
(1) $\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$
where $x_{i}$ denotes each instance in the training dataset and $p_{\theta}(x_{i})$ represents the model’s probability distribution parameterized by $\theta$. And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub (Husain et al., 2020; Kocetkov et al., 2022; Gao et al., 2020).
Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle (Wang et al., 2021; Fried et al., 2022; Nijkamp et al., 2023; Zheng et al., 2023) and also instruction fine-tuning (Luo et al., 2023; Rozière et al., 2023). And recent large language models pretrained on a mixture of code and natural language like ChatGPT (Schulman et al., 2022), GPT4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
2.2. Perturbation-Based Detection of Machine-Generated Text
In the realm of machine-generated text detection, perturbation based method like DetectGPT (Mitchell et al., 2023) stands as the state-of-the-art technology (Yang et al., 2023). In this section, we take DetectGPT as an example to illustrate the idea of perturbation-based detection methods. DetectGPT distinguishes between machine and human-generated text by analyzing the patterns in their probability scores (Mitchell et al., 2023). The core idea is that when a text $x$ generated from a machine is subtly changed to $\tilde{x}$ through a perturbation process $q(\cdot|x)$ (e.g., MLM with T5 (Ahmad et al., 2021)), there is a sharper decline in its log probability scores $\log p_{\theta}(x)$ than that in human-generated text. This is because a machine-generated text is usually more predictable and tightly bound to the patterns it was trained on, leading to a distinct negative curvature in log probability when the text is perturbed. By contrast, human-written texts are characterized by a rich diversity that reflects a blend of experiences and cognitive processes. As a result, it doesn’t follow such predictable patterns, and its log probability scores $\log p_{\theta}(x)$ do not plummet as dramatically when similarly perturbed. Based on such discrepancy, we can define a likelihood discrepancy score for each input code to measure the drop of log probability after perturbation.
(2) $\mathbf{d}\left(x,p_{\theta},q\right)\triangleq\log p_{\theta}(x)-\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log p_{\theta}(\tilde{x})$
By inspecting these scores, we can detect the source of $x$. A significant drop indicates machine authorship and a smaller change suggests a human creator. This method effectively captures the more nuanced and variable nature of human-generated text compared to the more formulaic and patterned output of language models.
3. Empirical Analysis
In this section, we conduct a comparative analysis of the distinct features of machine-generated and human-written code. We choose to assess code conciseness through length variability, evaluate vocabulary breadth via lexical diversity, understand common coding patterns by analyzing token frequency, identify stylistic differences through token category proportions, and gauge code fluency and predictability by measuring naturalness. These choices are aimed at providing a comprehensive analysis of coding styles between human and machine programmers.
3.1. Study Design
3.1.1. Length Variability
Code length stands as a cornerstone attribute when characterizing code (Barron et al., 1998; Albrecht and Gaffney, 1983; Rosenberg, 1997). The intricate balance of code length directly influences readability, maintainability, and even computational efficiency. A deeper grasp on how machine-generated code varies in length versus human-crafted code could reveal nuances about the inherent logic and preferences of LLMs, as well as the conventional coding patterns adopted by human developers. We investigate two metrics that characterize code length, namely, the number of tokens and the number of lines.
3.1.2. Lexical Diversity
Lexical diversity indicates the richness and variety of vocabulary present in a corpus. In the context of programming, this refers to the diversity in variable names, functions, classes, and reserved words. Analyzing lexical diversity offers a deeper understanding of the creativity, expressiveness, and potential complexity of code segments. There are two important empirical laws in both natural and programming languages revealing the lexical diversity: Zipf’s law (Zipf, 2016) and Heaps’ law (Heaps, 1978).
Zipf’s Law: It is initially identified in natural languages (Zipf, 2016), and later verified in the scope of programming languages (Zhang, 2008). It states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ (Zipf, 2016). In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed.
Heaps’ Law: Also verified in both text and code domain(Heaps, 1978; Zhang, 2009), Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human lexical usage.
3.1.3. Token Frequency
Token frequency stands for the occurrence of distinct tokens in the code corpus. The attribute indicates the core vocabulary utilized by human and machine programmers, shedding light on their coding preferences and tendencies.
3.1.4. Token Category Proportion
Token category proportion refers to the proportion of tokens belonging to different categories (e.g., keyword, identifiers) in the code corpus. Understanding the proportion of token categories in code is akin to dissecting the anatomy of a language. It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human and machine-authored code.
To delve into the token category proportion, we analyze code with tree-sitter [^1] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table 1. We then compute the proportion of each category in the code corpus.
Table 1. Studied categories of Python code tokens
Category Tree-sitter Types
keyword
def, return, else, if, for, while, $\ldots$
identifier
identifier, type_identifier
literal
string_content, integer, true, false, $\ldots$
operator
¿, ¡, =, ==, +, $\ldots$
syntactic symbol
comment
comment
whitespace
[…]
3.1.5. Naturalness
The concept of code naturalness suggests that programming languages share a similar degree of regularity and predictability with natural languages (Hindle et al., 2016). This idea has been operationalized by employing language models to assess the probability of a specific token’s occurrence within a given context. Under this framework, we inspect how “natural” machine-generated code is compared to human-written code. We employ two metrics, likelihood and rank to measure the naturalness of each token in the studied code corpus. The likelihood stands for the probability $p$ of a token $x$ under the model $p_{\theta}$, denoted as $p_{\theta}(x)$. The rank of a token $x$ is the position of $x$ in the sorted list of all tokens based on their likelihoods, denoted as $r_{\theta}(x)$. Both metrics evaluate how likely a token is preferred by the model (Ippolito et al., 2020; Gehrmann et al., 2019; Solaiman et al., 2019). We calculate log scores on each token and then take the average to represent the whole code snippet as advised in (Gehrmann et al., 2019). To pinpoint the code elements that most significantly affect the score discrepancies, we also present the mean scores on different token categories in Table 1 for comparison.
3.2. Experimental Setup
We choose the state-of-the-art CodeLlama model (Rozière et al., 2023) to generate code. Limited by our computational resources, we use the version with 7B parameters. As for the decoding strategies, we adopt the top-$p$ sampling method (Holtzman et al., 2020) with $p$=$0.95$ following (Chen et al., 2021). The temperature $T$ is an important parameter controlling the diversity of the generated code (Holtzman et al., 2020). Since current LLMs on code are usually evaluated across different decoding temperatures (Zheng et al., 2023; Nijkamp et al., 2022; Chen et al., 2021; Rozière et al., 2023), we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings. The maximum length of the generated code is set to 512 tokens. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
3.3. Dataset Preparation
To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus (Husain et al., 2020), a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase. We used function signatures and their accompanying comments as prompts for the model, and we collected the corresponding bodies of these functions to represent human-written code.
While acknowledging that current models, including CodeLlama and even ChatGPT (Liu et al., 2023), may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet (Yetiştiren et al., 2023; Liu et al., 2023), the choice of this dataset is deliberate and insightful. Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval (Chen et al., 2021) or MBPP (Austin et al., 2021) dataset, offers a more representative assessment. It allows us to analyze the differences between human and machine’s code when faced with broader, practical applications.
3.4. Results and Analysis
We present the results and analysis regarding each code attribute introduced in Section 3.1.
3.4.1. Length Variability
(a) Token length ($T=0.2$)
(b) Token length ($T=1.0$)
(c) Line numbers ($T=0.2$)
(d) Line numbers ($T=1.0$)
Figure 1. Distribution of length for generated code under two different temperatures
Figure 1 presents the distribution of code length under different settings. For the temperature setting of $0.2$, machine-authored code exhibits more conciseness, both in token and line numbers. As the temperature increases to $T=1.0$, we witness a convergence of distributions. The gap narrows, yet the machine’s preference for relatively concise code persists. This reveals that higher temperatures induce more exploratory generative behavior in the model, leading to diverse coding styles.
One could hypothesize several reasons for these observed patterns. The propensity for conciseness at lower temperatures may reflect LLM’s training data, where probably concise solutions were more prevalent or deemed more ‘correct’. On the flip side, human developers, often juggling multiple considerations like future code extensions, comments for peer developers, or even personal coding style, might craft lengthier solutions. Furthermore, the narrowing of disparities at higher temperatures can be attributed to the model’s increased willingness to explore varied coding styles. At higher temperatures, the LLM possibly mimics a broader spectrum of human coding patterns, capturing the essence of diverse coding habits and styles found in its training corpus.
Finding 1: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
(a) Zipf’s Law ($T=0.2$)
(b) Zipf’s Law ($T=1.0$)
(c) Heaps’ Law ($T=0.2$)
(d) Heaps’ Law ($T=1.0$)
Figure 2. Comparison of lexical diversity on machine- and human-authored code
3.4.2. Lexical Diversity
Figure 2 offers a comprehensive understanding of coding tendencies of both human and machine programmers. Starting with Zipf’s law, Figure 1(a) and 1(b) both delineate similar trends for human and machine programmers, corroborating the law’s applicability. And we can observe machine’s heightened proclivity towards tokens ranked between 10 and 100, especially at $T$=$0.2$. Turning our attention to Heaps’ law, the near-linear trends in Figure 1(c)-1(d) reaffirm the law’s validity. Also, there’s a noticeable shallowness in the slope for machine’s code at $T$=$0.2$ revealing machine’s decreased lexical diversity.
The obvious differences at $T$=$0.2$ can be ascribed to human creativity and variability. The varied approaches and methodologies humans employ can lead to a diversified token usage within this range. Another plausible interpretation can be its risk aversion. A machine, especially at lower temperatures, might be reverting to familiar patterns ensuring correctness in code generation. Additionally, certain patterns within the training data might have been overemphasized due to model overfitting, leading the machine to a skewed preference.
Finding 2: Machines demonstrate a preference for a limited spectrum of frequently-used tokens, whereas human code exhibits a richer diversity in token selection.
3.4.3. Token Frequency
Table 2. Top 50 tokens from human- and machine-authored code from CodeLlama on CodeSearchNet with $T=0.2$
Rank Human-Authored Tokens Machine-Authored Tokens
1–10 . ( ) = ’ , : self ” [ . - , ( ) self : ” ’ if
11–20 ] if return in for not None 0 1 == = return not [ def raise isinstance ] == path
21–30 else + is name { } path data raise - name 0 __class__ __name__ None os { / } %
31–40 try * os len format get and True value isinstance else TypeError str ‘ __init__ is ¿ // the
41–50 args key % np i x kwargs except False or in 1 ; value kwargs #include + __str__ for ValueError
Table 2 lists the top 50 tokens from human- and machine-authored code when $T$=0.2. Due to space limit, we omit the results when $T=1.0$, which has a similar result. From the results, we have several noteworthy observations:
Common Tokens: Human- and machine-authored code shares a commonality in their usage of certain tokens, including punctuation marks such as ‘.’, ‘(’, ‘)’, and structural keywords such as ‘if’, ‘return’, and ‘else’. This is because LLMs acquire foundational coding syntax after being trained on extensive human-written code corpora.
Error Handling: Tokens associated with error handling like ‘raise’ and ‘TypeError’, are more prevalent in machine-authored code. This difference implies that machine programmers emphasize robustness and exception handling more explicitly.
Programming Paradigms: Tokens indicating object-oriented programming like ‘self’ and ‘__init__’ are prominent in both human- and machine-authored code, which illustrates the model’s training alignment with this paradigm. However, machine-authored code appears to favor more boilerplate tokens like ‘__class__’ and also ‘__name__’, which could stem from its training on diverse object-oriented codebase.
Finding 3: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
3.4.4. Token Category Proportion
Figure 3. Proportions of token categories in the studied code corpus
The analysis of token category proportions, as visualized in Figure 3, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The ‘keyword’, ‘operator’, ‘syntactic symbols’, and ‘whitespace’ proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers:
Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machines favor reusing established identifiers over creating new ones, possibly a reflection of its programming paradigm influenced by its training data.
Literal: Machine-authored code consistently shows a marginally higher inclination towards literals. Across both temperatures, the machine exhibits an increase in the literal proportion as opposed to the human-written code. This indicates a potential proclivity of machine-authored code to process raw data more frequently. This could be attributed to the machine’s broader exposure to varied tasks involving data manipulations.
Comment: Machine-authored code has much more comments when $T$=1.0. This observation hints machine’s increased emphasis on code documentation and explanation with higher temperatures, when it becomes less deterministic and more exploratory.
Finding 4: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
3.4.5. Naturalness
Figure 4. Distribution of naturalness scores for human- and machine-authored code
Figure 4 shows that there is a great discrepancy of naturalness between machine- and human-authored code. Compared to human-authored code, the log likelihood scores of machine-authored code are mostly higher and the log rank scores are mostly lower, indicating that machine’s code is more “natural” than human-written code. Such observation in source code is consistent with the findings in natural language (Gehrmann et al., 2019; Ippolito et al., 2020; Mitchell et al., 2023; Solaiman et al., 2019).
Table 3. The naturalness of different token categories
Category Log Likelihood Log Rank
Machine Human $\Delta$ Machine Human $\Delta$
keyword -1.701 -2.128 0.428 0.837 1.053 0.217
identifier -0.459 -0.874 0.415 0.163 0.378 0.215
literal -0.506 -1.364 0.858 0.152 0.630 0.479
operator -0.938 -1.835 0.897 0.367 0.872 0.504
symbol -0.868 -1.639 0.771 0.321 0.781 0.460
comment -1.503 -3.028 1.525 0.608 1.610 1.002
whitespace -1.131 -2.740 1.609 0.429 1.441 1.012
ALL -1.658 -0.827 0.831 0.319 0.811 0.492
Table 3 summarises the comparison results in terms of each token category at $T=0.2$. An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens. This highlights a distinctive aspect of coding practices: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns. Humans, however, influenced by individual styles and practices, exhibit a wider variety in their use of whitespaces. The distinct patterns in machine-generated whitespaces, therefore, point to an inherent variation in coding style between machine and human.
Finding 5: Machine-authored code exhibits higher “naturalness” than human-authored code, and the disparity is more pronounced in tokens such as comments, and whitespaces, which are more reflective of individual coding styles.
4. Detecting Machine Generated Code
The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces. This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those structural tokens that best characterize the machine’s preference. Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-generated code. Our approach adapts the perturbation-based detection framework DetectGPT (Mitchell et al., 2023) to the code domain.
4.1. Problem Formulation
We formulate machine-generated code detection as a classification task that predicts whether a given code snippet $x$ is produced by a source model $p_{\theta}$. For this purpose, we transform $x$ to an equivalent form $\tilde{x}$ through a perturbation process $q(\cdot|x)$. We anticipate a sharper decline in its naturalness score if $x$ is written by an LLM. The key problems here are how to define the naturalness score and how to design the perturbation process. We introduce the naturalness score and the perturbation strategy $q(\cdot|x)$ in our approach in the following sections.
4.2. Measuring Naturalness
Previous methods usually use the log likelihood of tokens to measure the naturalness of machine-authored content (Hindle et al., 2016; Solaiman et al., 2019). However, the log rank of tokens shows better performance comparing the naturalness of machine- and human-authored text (Mitchell et al., 2023; Gehrmann et al., 2019), because it offers a smoother and robust representation of token preference.
Unlike DetectGPT which directly calculates the log likelihood of tokens, we adopt the Normalized Perturbed Log Rank (NPR) score (Su et al., 2023) to capture the naturalness. The NPR score is formally defined as:
(3) $\mathbf{NPR}\left(x,p_{\theta},q\right)\triangleq\frac{\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log r_{\theta}\left(\tilde{x}\right)}{\log r_{\theta}(x)},$
where $\log r_{\theta}(x)$ is the logarithm of the rank order of text $x$ sorted by likelihood under model $p_{\theta}$. In practice, $\mathbf{NPR}\left(x,p_{\theta},q\right)$ has been demonstrated to be more accurate for differentiating text origins, outperforming the log likelihood discrepancy $\mathbf{d}\left(x,p_{\theta},q\right)$ (Su et al., 2023).
4.3. Perturbation Strategy
Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in structural coding style. Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below. Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
Space Insertion
Let $C$ represent the set of all possible locations to insert spaces in a code segment. We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations. For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$. Mathematically, this can be represented as:
(4) $n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$
Newline Insertion
We split the codes into lines and obtain a set $L$ of lines. A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations. For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$. Formally, this is expressed as:
(5) $n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$
We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations. Through this step, we instill randomness at a granular structural level, thereby amplifying the perturbation’s efficacy. Our perturbation strategy introduces several distinct advantages over the conventional methods (Mitchell et al., 2023; Su et al., 2023) using MLM to perturb the code, which will be discussed in Section 6.2.
Data: code $x$, source model $\mathcal{M}$, number of perturbations $k$, decision threshold $\epsilon$, parameters $\alpha$, $\beta$, $\lambda_{\text{spaces}}$, and $\lambda_{\text{newlines}}$
// Detection using NPR score on Perturbed Samples
1 for $i\leftarrow 1$ to $k$ do
       // Random decision for type of perturbation
2       $p\sim\mathcal{U}(0,1)$;
3       if $p\leq 0.5$ then
             // Spaces Insertion for Perturbed Sample
4             Let $C$ represent the set of all possible locations to insert spaces in $x$;
5             Select $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$;
6             for each location $c\in C_{s}$ do
7                   $n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}})$;
8                   Insert $n_{\text{spaces}}(c)$ spaces at location $c$ in $x$;
9
10             end for
11
12      else
             // Newlines Insertion for Perturbed Sample
13             Split the perturbed code $x$ into a set $L$ of lines;
14             Select $L_{n}\subseteq L$ such that $|L_{n}|=\beta\times|L|$;
15             for each line $l\in L_{n}$ do
16                   $n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}})$;
17                   Insert $n_{\text{newlines}}(l)$ newlines after line $l$ in $x$;
18
19             end for
20
21       end if
22      Store the perturbed code as $\tilde{x}_{i}$;
23
24 end for
25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\mathbf{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\mathbf{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
27 Calculate variance: $\tilde{\sigma}_{x}^{2}\leftarrow\frac{1}{k-1}\sum_{i}\left(\mathbf{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)-\tilde{\mu}\right)^{2}$;
28 if $\frac{\text{NPR}_{x}}{\sqrt{\tilde{\sigma}_{x}}}>\epsilon$ then
       return true ;
        // Probably machine-generated
29
30else
       return false ;
        // Probably human-written
31
32 end if
Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Structured Code Perturbation
Algorithm 1 summarizes the entire workflow of DetectCodeGPT. Our algorithm harnesses structured code perturbation to differentiate between human- and machine-authored code. At the core of our approach is the strategic insertion of spaces (Lines 5-9) and newlines (Lines 11-16) in code, a process that simulates the inherent randomness in human coding habits. The algorithm operates by generating perturbed versions of the code and then evaluating their NPR scores (Lines 20-22) with respect to the source model $\mathcal{M}$.
The threshold parameter $\epsilon$ in Line 23, pivotal for making the detection decision, offers flexibility in catering to different application scenarios. By adjusting $\epsilon$, users can balance between false positives and false negatives, tailoring the detection sensitivity according to the specific needs of the deployment context.
5. Evaluation
We conduct experiments to evaluate the effectiveness of DetectCodeGPT, aiming to answer the following research questions.
  • RQ1: How effectively does our method distinguish between machine-generated and human-written code?
  • RQ2: To what extent do individual components influence the overall performance of our method?
  • RQ3: What is the impact of varying the number of perturbations on the detection performance?
  • RQ4: How resilient is our method to cross-model code source detection?
5.1. Datasets
Apart from CodeSearchNet (Husain et al., 2020) described in Section 3.3, we select Python code from The Stack (Kocetkov et al., 2022) as another evaluation dataset. Similar to CodeSearchNet, The Stack provides codes from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version (Manh et al., 2023) of this dataset and also concatenate the function definitions with their corresponding comments as prompts. For each combination of dataset and model, we construct 500 human and machine code pairs for evaluation. The maximum length of both human- and machine-authored code is trimmed to 128 tokens.
5.2. Evaluation Metric
Following prior works (Mitchell et al., 2023; Su et al., 2023), our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code. Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as:
[…]
where $t$ denotes varying threshold values. It provides a comprehensive view of performance across all possible thresholds, making it threshold-independent. This makes the metric both interpretable and insightful, offering a clearer picture of the model’s discriminating capabilities.
5.3. Baselines and Models
Our evaluation is benchmarked against a diverse range of zero-shot machine-generated text detection techniques, all of which leverage the predicted token-wise conditional distributions of the source model for detection. A supervised baseline is also included to demonstrate the effectiveness of our method.
  • Log $p(x)$ (Solaiman et al., 2019): Utilizes the source model’s average token-wise log probability to gauge code naturalness. Machine-generated code tends to have a higher score.
  • Entropy (Gehrmann et al., 2019): Interprets high average entropy in the model’s predictive distribution as indicative of machine generation.
  • (Log-) Rank (Gehrmann et al., 2019; Ippolito et al., 2020): The average observed rank or log rank of each token in the LLM prediction, with machine-generated passages typically showing smaller average values.
  • DetectGPT (Mitchell et al., 2023): Leverages the log probability of the original code and its perturbed variants to compute the perturbation discrepancy gap.
  • DetectLLM (Su et al., 2023): Introduces two methods, one blends log likelihood with log rank to compute LRR, and the other improves DetectGPT by incorporating the NPR score.
  • Supervised Method: a simple supervised detection method that predicts the authorship of a given code snippet using a binary classifier. Following OpenAI’s RoBERTa-based (Liu et al., 2019) GPT-2 detector [^2] https://github.com/openai/gpt-2-output-dataset/tree/master/detector, we employ CodeBERT (Feng et al., 2020) with an MLP head as the classifier and train it on a combination of 1000 samples generated by each model at each setting.
As for models to generate machine’s code, in consideration of GPU memory limitations, we judiciously selected a range of advanced LLMs with varying parameter counts, from 1 billion to 7 billion. We selected Incoder (Fried et al., 2022), Phi-1 (Gunasekar et al., 2023), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023), CodeGen2 (Nijkamp et al., 2023) and CodeLlama (Rozière et al., 2023) from Huggingface  [^3] https://huggingface.co/models. This diverse array of models, each trained on distinct datasets and boasting different capacities, offers a comprehensive spectrum for evaluating our method.
5.4. Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy (Holtzman et al., 2020) with $p$=$0.95$ following (Chen et al., 2021). We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section 3.2. The maximum length constraint for generated codes was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model (Allal et al., 2023), revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}=3$ and $\lambda_{\text{newlines}}=2$. For all experiments, we maintained a consistent configuration of generating $50$ perturbations. For the DetectGPT and DetectLLM tasks, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model (Wang et al., 2023). All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
Table 4. Performance (AUROC) of various detection methods
Dataset Code LLM Detection Methods
$\log p(x)$ Entropy Rank Log Rank  DetectGPT LRR NPR  Supervised  DetectCodeGPT
CodeSearchNet ($T=0.2$) Incoder (1.3B) 0.9810 0.1102 0.8701 0.9892 0.4735 0.9693 0.8143 0.9426 0.9896
Phi-1 (1.3B) 0.7881 0.4114 0.6409 0.7513 0.7210 0.4020 0.7566 0.3855 0.8287
StarCoder (3B) 0.9105 0.2942 0.7585 0.9340 0.6949 0.9245 0.9015 0.7712 0.9438
WizardCoder (3B) 0.9079 0.2930 0.7556 0.9120 0.6450 0.7975 0.8677 0.7433 0.9345
CodeGen2 (3.7B) 0.7028 0.4411 0.7328 0.7199 0.6051 0.7997 0.6177 0.5327 0.8802
CodeLlama (7B) 0.8850 0.3174 0.7265 0.9016 0.8212 0.8332 0.5890 0.7496 0.9095
CodeSearchNet ($T=1.0$) Incoder (1.3B) 0.7724 0.4167 0.7797 0.7876 0.6258 0.7427 0.6801 0.6761 0.7882
Phi-1 (1.3B) 0.6118 0.4588 0.5709 0.6299 0.7492 0.4528 0.7912 0.4158 0.8365
StarCoder (3B) 0.6574 0.4844 0.6987 0.6822 0.6505 0.7050 0.6751 0.6299 0.6918
WizardCoder (3B) 0.8319 0.3363 0.7273 0.8338 0.5972 0.6965 0.7516 0.7068 0.8392
CodeGen2 (3.7B) 0.4484 0.6263 0.6584 0.4632 0.4797 0.5530 0.5208 0.4024 0.6798
CodeLlama (7B) 0.6463 0.4855 0.6759 0.6656 0.6423 0.6768 0.6515 0.6442 0.7239
The Stack ($T=0.2$) Incoder (1.3B) 0.9693 0.1516 0.8747 0.9712 0.6061 0.9638 0.8571 0.9291 0.9727
Phi-1 (1.3B) 0.8050 0.4318 0.6766 0.7622 0.7295 0.4022 0.8106 0.4640 0.8578
StarCoder (3B) 0.9098 0.3077 0.7843 0.9329 0.6824 0.9135 0.9233 0.7715 0.9274
WizardCoder (3B) 0.9026 0.3196 0.7963 0.9010 0.6385 0.7742 0.8574 0.7794 0.9243
CodeGen2 (3.7B) 0.7171 0.4051 0.7930 0.7301 0.5288 0.7604 0.5670 0.4520 0.8513
CodeLlama (7B) 0.8576 0.3565 0.7366 0.8793 0.8087 0.8358 0.5436 0.7619 0.8852
The Stack ($T=1.0$) Incoder (1.3B) 0.7310 0.4591 0.7673 0.7555 0.6124 0.7446 0.6787 0.6846 0.7833
Phi-1 (1.3B) 0.7841 0.4205 0.6666 0.7475 0.6718 0.4106 0.7755 0.4984 0.8376
StarCoder (3B) 0.6333 0.5025 0.7010 0.6609 0.5896 0.7080 0.6638 0.7243 0.6890
WizardCoder (3B) 0.8293 0.3459 0.7484 0.8223 0.6377 0.6436 0.7929 0.7766 0.8384
CodeGen2 (3.7B) 0.4816 0.6046 0.5631 0.4956 0.4337 0.5740 0.5178 0.4265 0.6660
CodeLlama (7B) 0.5929 0.5260 0.6451 0.6091 0.6116 0.6365 0.6226 0.7494 0.6595
5.5. Detection Performance (RQ2)
Table 4 delineates the results of various methods. According to the results, DetectCodeGPT consistently outperforms baseline methods. In an impressive 21 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama. The high AUROC values achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
We can observe that the challenge of detection notably increases at a temperature setting of $T$=$1.0$ than $T$=$0.2$. This is possibly due to the higher randomness at this temperature, where models are likely to generate outputs with greater diversity. Despite these increased difficulties, the proposed method maintains its leading position in detection accuracy.
It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised approach. The unstable performance of the supervised model reveals that it’s hard to detect machine-generated codes relying on the training data. The ability of our zero-shot approach to adapt and retain high performance under different levels of generative temperature variations is indicative of its sophisticated detection capability, making it a valuable tool for maintaining the integrity of codebases against the infiltration of machine-generated code in practical applications.
5.6. Ablation Study (RQ3)
Table 5. Performance of different perturbation strategies
Perturb. Type MLM   Newline Space   Newline&Space
$T=0.2$ 0.5436 0.8703 0.8639 0.8852
$T=1.0$ 0.6226 0.6453 0.6504 0.6660
In our ablation study, we compared the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset. The results summarized in Table 5 illuminates the comparative advantage of our structured perturbation approach. We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM (Mitchell et al., 2023; Su et al., 2023) for natural language. Also, the combination of newline and space perturbations further enhances the detection performance, with the highest AUROC score of 0.8852 at $T=0.2$ and 0.6660 at $T=1.0$. The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
5.7. Impact of Perturbation Count (RQ4)
To gauge the impact of perturbation count on the efficacy of our method, we conducted experiments with varying numbers of perturbations.
Table 6. Impact of varying the number of perturbations
#Perturbations 10 20 50 100 200
$T=0.2$ 0.6537 0.8825 0.8852 0.8855 0.8846
$T=1.0$ 0.5558 0.6584 0.6660 0.6660 0.6662
Results in Table 6 reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach. Notably, an increase to 20 perturbations already yields robust detection performance, with further increments leading to diminishing improvements. This suggests that our method requires a relatively small number of perturbations to effectively discern between human- and machine-authored code. This implies that our method is not only effective but also resource-efficient.
5.8. Performance of Cross-Model Detection (RQ5)
Figure 5. Cross-model detection performance
In previous sections, we primarily assessed the efficacy of DetectCodeGPT within a white-box framework, where the detection model can be the same as the source model. Nonetheless, in real-world applications, it is often infeasible to access to the original code generation model for detection. To simulate such conditions, we engaged in cross-model detection experiments, wherein distinct models were utilized to compute the naturalness scores for detecting code generated by others. These evaluations were carried out using The Stack dataset at a temperature of 0.2.
The results presented in Figure 5 highlight DetectCodeGPT’s adaptability in cross-model detection. While the algorithm excels in the white-box setting, its performance endures with only a slight reduction when subjected to cross-model application. For instance, StarCoder, when detecting code generated by WizardCoder and CodeLlama, yields AUROC scores of 0.92 and 0.87, respectively, compared to an AUROC of 0.93 when detecting its own output. And we can also notice a performance decrease when detecting CodeGen2’s output. This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text (Nijkamp et al., 2023). However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in (Mireshghallah et al., 2023).
These results indicate that DetectCodeGPT is a model-free method that is robust against model discrepancies.
6. Discussion
In this section, we will analyze why our method is effective and discuss its strengths and limitations.
6.1. Why is DetectCodeGPT Effective?
We attribute the effectiveness of our DetectCodeGPT to the following two factors:
Preservation of Code Correctness
While the perturbation method in DetectGPT masks continuous spans of text and requires another LLM to recover, such recovery often leads to structurally compromised code. The unique syntax and constraints inherent to programming languages mean that even minor misinterpretations (e.g., misuse of an identifier) can render the code non-functional. The originally masked spans could be the only correct solution to recover the code and the recovered version from LLM bring minor mistakes easily. Such code-cracking perturbations will possibly lead to both a decrease in naturalness scores for human’s and machine’s codes, violating the assumption that the perturbation will have minimal impact on the code’s naturalness score if it is human-written in Section 2.2. In contrast, inserting newlines and spaces does not affect the correctness of the code in most cases. So our perturbation strategy ensures that the underlying structure and correctness of the code remain intact, thereby ensuring the effectiveness of our method.
Emulation of Human Randomness
As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of spaces and newlines than machines during coding. For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to structure the code in a more standardized and compact manner. Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section 3.
6.2. Strength of DetectCodeGPT
DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding preferences. We summarize the strengths of DetectCodeGPT as follows:
Enhanced Efficiency
Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection (Mitchell et al., 2023; Su et al., 2023), as verified in Section 5.7, perturbing code using MLM will lead to huge computational cost. On the contrary, our method will save much more time and bring more efficiency. This translates to a significant conservation of computational resources, thus expediting the detection process and enhancing scalability. Mathematically, if $T_{\text{DetectGPT}}$ and $T_{\text{Ours}}$ denote the computational times of DetectGPT and our method respectively, we know that $T_{\text{Ours}}\ll T_{\text{DetectGPT}}$.
Superior Zero-Shot Performance
DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets. This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated in Section 5.8. In comparison to the supervised methods, which may suffer from data drift or the need for constant updates, our approach is not only more adaptable but also delivers better performance across various data sources. This makes DetectCodeGPT an agile and reliable tool for machine-generated code detection in a variety of contexts.
6.3. Limitations and Future Directions
The main limitations of our work lie in the following two aspects: Firstly, due to the time and computational constraints, we only focus on a set of LLMs within 7B parameters for both empirical analysis and the machine-generated code detection task. As the landscape of LLMs rapidly evolves, incorporating a wider array of more and larger LLMs could significantly bolster the generalizability and robustness of our findings.
Secondly, our current analysis centers exclusively on Python code, while the features of other programming languages may not be fully explored and the effectiveness of our method on other languages is not verified. However, based on the analysis of our method in Section 6.2, we believe that our method can be effectively generalized to other programming languages, especially where the functionality of code won’t be much affected after inserting newlines and spaces like C/C++, Java, and JavaScript.
Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness, such as $T$=$1.0$. We note from Table 4 that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement. Although ensembling multiple detection models may help to improve the detection performance (Mireshghallah et al., 2023), we look forward to explore more effective perturbation strategies based on code structure to further enhance the detection efficacy.
Additionally, the development of watermarking strategies (Sun et al., 2023) for machine-generated code represents an intriguing avenue for future research. Given that our method can effectively identify machine’s code, it could be adapted to assist in the creation of watermarking techniques. For instance, embedding a watermark could be achieved by inserting newlines and spaces in a code-specific pattern, which would increase detectablility without compromising functionality.
Moreover, our work also highlights the need for the development of more robust detection methods against possible attacks (Hu et al., 2023; Krishna et al., 2023). As discussed in Section 3, an important difference between machine- and human-authored code is naturalness. We can possibly leverage the insertion of whitespace to imitate a less ‘natural’ code style and thereby make machine’s code more difficult to detect. This potential strategy highlights the need for the development of more robust detection methods against possible attacks (Hu et al., 2023; Krishna et al., 2023).
7. Related Work
Recently, there has been much effort in detecting machine-generated text (Yang et al., 2023; Wu et al., 2023b). The two main categories of detection methods are zero-shot and training-based methods, and our DetectCodeGPT falls into the former category.
As for zero-shot methods, they are usually based on the discrepancy between likelihood and rank information of human and machine’s texts (Gehrmann et al., 2019; Ippolito et al., 2020; Solaiman et al., 2019). Leveraging the hypothesis in DetectGPT (Mitchell et al., 2023) that machine-generated text often has a negative curvature in the log probability when the text is perturbed, many perturbation based methods have been proposed (Mitchell et al., 2023; Su et al., 2023; Bao et al., 2023). These methods usually perturb the text by masking a continuous span of tokens and then recover the perturbed text using another LLM like T5 (Ahmad et al., 2021). The benefit of these methods is that they are zero-shot and can be applied to any LLM without access to training data. However, the perturbation process is time-consuming and computationally expensive,
When it comes to the training-based methods, fine-tuning the RoBERTa(Liu et al., 2019) or T5 (Ahmad et al., 2021) model with data collected from different model families at different decoding settings is a common practice (Tian et al., 2023; Zhan et al., 2023; Chen et al., 2023). Additional information like graph structure (Liu et al., 2022), perplexity from proxy models (Wu et al., 2023a) have been shown to be helpful for detection. Moreover, techniques like adversarial training (Hu et al., 2023) and contrastive learning (Liu et al., 2022) have also been proposed to improve the detection performance. The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model.
However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text (Lee et al., 2023). The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation (Lee et al., 2023; Sun et al., 2023). The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism (Cox et al., 2002; Collberg and Thomborson, 1999). Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
8. Conclusion
In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the age of AI. By pioneering a comparative analysis across five different attributes, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces. Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the structural nuances of code with a novel perturbation strategy that is simple yet effective. The empirical success of DetectCodeGPT against several benchmarks confirms its potential to bolster security measures in programming and establish a more symbiotic relationship between human developers and LLMs. Our work marks a significant step towards responsible AI use in software engineering, ensuring clarity in code authorship and maintaining the integrity of code as a craft.
Data Availability
All the experimental data and source code used in this work are available at https://github.com/YerbaPage/DetectCodeGPT.

### v2

@@
-Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers
-Abstract.
+Index Terms: Machine-generated code detection, large language models, code generation, empirical analysis
+Abstract
 Large language models have catalyzed an unprecedented wave in code generation.
@@
 In this paper, we carefully study the specific patterns that characterize machine and human-authored code.
-Through a rigorous analysis of code attributes such as length, lexical diversity, and naturalness, we expose unique patterns inherent to each source.
-We particularly notice that the structural segmentation of code is a critical factor in identifying its provenance.
-Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct structural patterns of code.
+Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source.
+We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance.
+Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct stylized patterns of code.
 Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency.
 Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
-copyright: none
-1. Introduction
-The advent of large language models (LLMs) such as Codex (Chen et al., 2021) and ChatGPT (OpenAI, 2022) has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect.
-LLMs trained on code (Allal et al., 2023; Nijkamp et al., 2023; Zheng et al., 2023; Rozière et al., 2023), sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation (Yetiştiren et al., 2023; Xu et al., 2022).
+Index Terms:
+Machine-generated code detection, large language models, code generation, empirical analysis
+1 Introduction
+The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect.
+LLMs trained on code [3, 4, 5, 6], sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation [7, 8].
 Their ability to generate syntactically correct and functionally robust code challenges traditional dynamics and introduces a new era of efficiency and innovation in software creation, maintenance, and evolution.
@@
 Addressing these concerns is pivotal in maintaining a transparent and secure software development lifecycle.
-Recently, there has been a growing research trend in detecting machine-generated texts (Yang et al., 2023; Zhao et al., 2023).
-Perturbation-based methods like DetectGPT (Mitchell et al., 2023), have achieved state-of-the-art results in identifying machine-generated text.
+Recently, there has been a growing research trend in detecting machine-generated texts [9, 10].
+Perturbation-based methods like DetectGPT [11], have achieved state-of-the-art results in identifying machine-generated text.
 DetectGPT employs likelihood score discrepancies between the original text and its perturbed variants to enhance detection accuracy.
-And the perturbation process is performed by a pretrained language model (e.g., T5 (Raffel et al., 2020)) to mask portions of the original text and then reconstruct it.
-While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid structural constraints.
-Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code (Lee et al., 2023).
-In this paper, we conduct a comparative analysis of the patterns between machine and human-written code from five aspects, including length, lexical diversity, token frequency, token category proportions, and naturalness.
-Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces.
+And the perturbation process is performed by a pretrained language model (e.g., T5 [12]) to mask portions of the original text and then reconstruct it.
+While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid syntactic constraints.
+Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code [13].
+In this paper, we conduct a comparative analysis of the patterns between machine- and human-authored code from three aspects, including lexical diversity, conciseness, and naturalness.
+Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespace tokens.
 Capitalizing on our findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code.
-Our DetectCodeGPT improves DetectGPT by capturing the distinct structural patterns of code.
+Our DetectCodeGPT improves DetectGPT by capturing the distinct styles of code.
 Diverging from conventional techniques that depend on external pretrained model (e.g.
-CodeT5 (Wang et al., 2021)) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code corpus by strategically inserting spaces and newlines.
-Eschewing the need for an LLM for perturbations, our method provides a much higher efficiency, Since the proposed perturbation strategy well captures the structural distinctions between machine and human code, our DetectCodeGPT also ensures higher efficacy.
-To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models (LLMs), ranging from 1.3 to 7 billion parameters, and operating at two different decoding temperatures.
-The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward (Yang et al., 2023; Tian et al., 2023), detecting machine-authored code presents a unique challenge.
-Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement with an average absolute increase of 6.46% in AUROC.
+CodeT5 [14]) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code by strategically inserting spaces and newlines.
+Eschewing the need for an LLM for perturbations, our method provides much higher efficiency.
+Since the proposed perturbation strategy well captures the stylistic distinctions between machine- and human-authored code, our DetectCodeGPT also ensures higher efficacy.
+To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models, ranging from 1.3 to 7 billion parameters, and operating at two different decoding conditions.
+The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward [9, 15], detecting machine-authored code presents a unique challenge.
+Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement.
 Our contributions can be summarized as follows:
@@
 Our study sheds light on essential insights that further advance the utility of LLMs in coding.
-  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct structural patterns of code.
+  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct stylistic patterns of code.
   • We extensively evaluate the DetectCodeGPT across a variety of settings and show the effectiveness of our approach.
-2. Background
-2.1. Large Language Models for Code
-GPT language models (Radford et al., 2018, 2019) based on Transformer (Vaswani et al., 2017) decoder has achieved remarkable success in natural language processing tasks (Ray, 2023).
-In the file of code generation, Codex (Chen et al., 2021) and AlphaCode (Li et al., 2022) are pioneering works to train large language models on code.
-These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens (Radford et al., 2018).
+2 Background
+2.1 Large Language Models for Code
+GPT language models [16, 17] based on Transformer [18] decoder has achieved remarkable success in natural language processing tasks [19].
+In the file of code generation, Codex [1] and AlphaCode [20] are pioneering works to train large language models on code.
+These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens [16].
 Mathematically, this is captured as:
-(1) $\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$
+$\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$ (1)
 where $x_{i}$ denotes each instance in the training dataset and $p_{\theta}(x_{i})$ represents the model’s probability distribution parameterized by $\theta$.
-And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub (Husain et al., 2020; Kocetkov et al., 2022; Gao et al., 2020).
-Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle (Wang et al., 2021; Fried et al., 2022; Nijkamp et al., 2023; Zheng et al., 2023) and also instruction fine-tuning (Luo et al., 2023; Rozière et al., 2023).
-And recent large language models pretrained on a mixture of code and natural language like ChatGPT (Schulman et al., 2022), GPT4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
-2.2. Perturbation-Based Detection of Machine-Generated Text
-In the realm of machine-generated text detection, perturbation based method like DetectGPT (Mitchell et al., 2023) stands as the state-of-the-art technology (Yang et al., 2023).
+And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub [21, 22, 23].
+Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle [14, 24, 4, 5] and also instruction fine-tuning [25, 6].
+And recent large language models pretrained on a mixture of code and natural language like ChatGPT [26], GPT4 [27] and LLaMA [28] has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
+2.2 Perturbation-Based Detection of Machine-Generated Text
+In the realm of machine-generated text detection, perturbation based method like DetectGPT [11] stands as the state-of-the-art technology [9].
 In this section, we take DetectGPT as an example to illustrate the idea of perturbation-based detection methods.
-DetectGPT distinguishes between machine and human-generated text by analyzing the patterns in their probability scores (Mitchell et al., 2023).
-The core idea is that when a text $x$ generated from a machine is subtly changed to $\tilde{x}$ through a perturbation process $q(\cdot|x)$ (e.g., MLM with T5 (Ahmad et al., 2021)), there is a sharper decline in its log probability scores $\log p_{\theta}(x)$ than that in human-generated text.
+DetectGPT distinguishes between machine and human-generated text by analyzing the patterns in their probability scores [11].
+The core idea is that when a text $x$ generated from a machine is subtly changed to $\tilde{x}$ through a perturbation process $q(\cdot|x)$ (e.g., MLM with T5 [29]), there is a sharper decline in its log probability scores $\log p_{\theta}(x)$ than that in human-generated text.
 This is because a machine-generated text is usually more predictable and tightly bound to the patterns it was trained on, leading to a distinct negative curvature in log probability when the text is perturbed.
@@
 Based on such discrepancy, we can define a likelihood discrepancy score for each input code to measure the drop of log probability after perturbation.
-(2) $\mathbf{d}\left(x,p_{\theta},q\right)\triangleq\log p_{\theta}(x)-\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log p_{\theta}(\tilde{x})$
+$\mathbf{d}\left(x,p_{\theta},q\right)\triangleq\log p_{\theta}(x)-\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log p_{\theta}(\tilde{x})$ (2)
 By inspecting these scores, we can detect the source of $x$.
@@
 This method effectively captures the more nuanced and variable nature of human-generated text compared to the more formulaic and patterned output of language models.
-3. Empirical Analysis
-In this section, we conduct a comparative analysis of the distinct features of machine-generated and human-written code.
-We choose to assess code conciseness through length variability, evaluate vocabulary breadth via lexical diversity, understand common coding patterns by analyzing token frequency, identify stylistic differences through token category proportions, and gauge code fluency and predictability by measuring naturalness.
-These choices are aimed at providing a comprehensive analysis of coding styles between human and machine programmers.
-3.1. Study Design
-3.1.1. Length Variability
-Code length stands as a cornerstone attribute when characterizing code (Barron et al., 1998; Albrecht and Gaffney, 1983; Rosenberg, 1997).
-The intricate balance of code length directly influences readability, maintainability, and even computational efficiency.
-A deeper grasp on how machine-generated code varies in length versus human-crafted code could reveal nuances about the inherent logic and preferences of LLMs, as well as the conventional coding patterns adopted by human developers.
-We investigate two metrics that characterize code length, namely, the number of tokens and the number of lines.
-3.1.2. Lexical Diversity
+3 Empirical Analysis
+In this section, we conduct a comparative analysis of the distinct features of machine- and human-authored code.
+3.1 Study Design
+To gain insights into distinctions between human and machine programmers, we consider three primary aspects that are relevant to coding styles, namely diversity, conciseness, and naturalness [30, 31, 32, 33, 34], which can be measured by specific metrics.
+3.1.1 Lexical Diversity
 Lexical diversity indicates the richness and variety of vocabulary present in a corpus.
@@
 Analyzing lexical diversity offers a deeper understanding of the creativity, expressiveness, and potential complexity of code segments.
-There are two important empirical laws in both natural and programming languages revealing the lexical diversity: Zipf’s law (Zipf, 2016) and Heaps’ law (Heaps, 1978).
-Zipf’s Law: It is initially identified in natural languages (Zipf, 2016), and later verified in the scope of programming languages (Zhang, 2008).
-It states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ (Zipf, 2016).
-In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed.
-Heaps’ Law: Also verified in both text and code domain(Heaps, 1978; Zhang, 2009), Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
-We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human lexical usage.
-3.1.3. Token Frequency
+There are four important empirical metrics in both natural and programming languages revealing the lexical diversity: token frequency, syntax element distribution, Zipf’s law [35] and Heaps’ law [36].
 Token Frequency stands for the occurrence of distinct tokens in the code corpus.
 The attribute indicates the core vocabulary utilized by human and machine programmers, shedding light on their coding preferences and tendencies.
-3.1.4. Token Category Proportion
-Token category proportion refers to the proportion of tokens belonging to different categories (e.g., keyword, identifiers) in the code corpus.
-Understanding the proportion of token categories in code is akin to dissecting the anatomy of a language.
-It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human and machine-authored code.
-To delve into the token category proportion, we analyze code with tree-sitter [^1] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table 1.
+Syntax Element Distribution refers to the proportion of syntax elements (e.g., keywords, identifiers) in the code corpus.
+Understanding the distribution of syntax elements in code is akin to dissecting the anatomy of a language.
+It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human- and machine-authored code.
+To delve into the syntax element distribution, we analyze code with tree-sitter [^1] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table I.
 We then compute the proportion of each category in the code corpus.
-Table 1. Studied categories of Python code tokens
+Zipf’s and Heaps’ Laws were initially identified in natural languages [35, 36], and later verified in the scope of programming languages [30, 31].
+Zipf’s law states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ [35].
+In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed.
+Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
+We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human lexical usage.
+Table I: Studied categories of Python code tokens
 Category Tree-sitter Types
@@
 operator
-¿, ¡, =, ==, +, $\ldots$
+<, >, =, ==, +, $\ldots$
 syntactic symbol
@@
 […]
-3.1.5. Naturalness
-The concept of code naturalness suggests that programming languages share a similar degree of regularity and predictability with natural languages (Hindle et al., 2016).
+3.1.2 Conciseness
+Conciseness stands as a cornerstone attribute when characterizing code [37, 33, 34].
+The intricate balance of code conciseness directly influences readability, maintainability, and even computational efficiency.
+We investigate two metrics that characterize code conciseness, namely, the number of tokens and the number of lines.
+Number of tokens gives us an indication of verbosity and complexity, showing the detailed composition of the code [37].
+Number of lines helps us understand organizational choices, as spreading code across more lines can reflect a focus on readability and structure [34].
+3.1.3 Naturalness
+The concept of code naturalness suggests that programming languages share a similar degree of regularity and predictability with natural languages [32].
 This idea has been operationalized by employing language models to assess the probability of a specific token’s occurrence within a given context.
 Under this framework, we inspect how “natural” machine-generated code is compared to human-written code.
-We employ two metrics, likelihood and rank to measure the naturalness of each token in the studied code corpus.
-The likelihood stands for the probability $p$ of a token $x$ under the model $p_{\theta}$, denoted as $p_{\theta}(x)$.
+Token Likelihood and Rank are two metrics that measure the naturalness of each token in the studied code corpus.
+The token likelihood stands for the probability $p$ of a token $x$ under the model $p_{\theta}$, denoted as $p_{\theta}(x)$.
 The rank of a token $x$ is the position of $x$ in the sorted list of all tokens based on their likelihoods, denoted as $r_{\theta}(x)$.
-Both metrics evaluate how likely a token is preferred by the model (Ippolito et al., 2020; Gehrmann et al., 2019; Solaiman et al., 2019).
-We calculate log scores on each token and then take the average to represent the whole code snippet as advised in (Gehrmann et al., 2019).
-To pinpoint the code elements that most significantly affect the score discrepancies, we also present the mean scores on different token categories in Table 1 for comparison.
-3.2. Experimental Setup
-We choose the state-of-the-art CodeLlama model (Rozière et al., 2023) to generate code.
+Both metrics evaluate how likely a token is preferred by the model [38, 39, 40].
+We calculate log scores on each token and then take the average to represent the whole code snippet as advised in [39].
+To pinpoint the code elements that most significantly affect the score discrepancies, we also present the mean scores on different syntax element categories in Table I for comparison.
+3.2 Experimental Setup
+We choose the state-of-the-art CodeLlama model [6] to generate code.
 Limited by our computational resources, we use the version with 7B parameters.
-As for the decoding strategies, we adopt the top-$p$ sampling method (Holtzman et al., 2020) with $p$=$0.95$ following (Chen et al., 2021).
-The temperature $T$ is an important parameter controlling the diversity of the generated code (Holtzman et al., 2020).
-Since current LLMs on code are usually evaluated across different decoding temperatures (Zheng et al., 2023; Nijkamp et al., 2022; Chen et al., 2021; Rozière et al., 2023), we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings.
+As for the decoding strategies, we adopt the top-$p$ sampling method [41] with $p$=$0.95$ following [1].
+The temperature $T$ is an important parameter controlling the diversity of the generated code [41].
+Since current LLMs on code are usually evaluated across different decoding temperatures [5, 42, 1, 6], we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings.
 The maximum length of the generated code is set to 512 tokens.
 All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
-3.3. Dataset Preparation
-To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus (Husain et al., 2020), a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase.
-We used function signatures and their accompanying comments as prompts for the model, and we collected the corresponding bodies of these functions to represent human-written code.
-While acknowledging that current models, including CodeLlama and even ChatGPT (Liu et al., 2023), may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet (Yetiştiren et al., 2023; Liu et al., 2023), the choice of this dataset is deliberate and insightful.
-Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval (Chen et al., 2021) or MBPP (Austin et al., 2021) dataset, offers a more representative assessment.
+3.3 Dataset Preparation
+To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus [21], a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase.
+We used function signatures and their accompanying comments as prompts for the model as in [1], and we collected the corresponding bodies of these functions to represent human-written code.
+While acknowledging that current models, including CodeLlama and even ChatGPT [43], may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet [7, 43], the choice of this dataset is deliberate and insightful.
+Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval [1] or MBPP [44] dataset, offers a more representative assessment.
 It allows us to analyze the differences between human and machine’s code when faced with broader, practical applications.
-3.4. Results and Analysis
+3.4 Results and Analysis
 We present the results and analysis regarding each code attribute introduced in Section 3.1.
-3.4.1. Length Variability
-(a) Token length ($T=0.2$)
-(b) Token length ($T=1.0$)
-(c) Line numbers ($T=0.2$)
-(d) Line numbers ($T=1.0$)
-Figure 1. Distribution of length for generated code under two different temperatures
-Figure 1 presents the distribution of code length under different settings.
-For the temperature setting of $0.2$, machine-authored code exhibits more conciseness, both in token and line numbers.
-As the temperature increases to $T=1.0$, we witness a convergence of distributions.
-The gap narrows, yet the machine’s preference for relatively concise code persists.
-This reveals that higher temperatures induce more exploratory generative behavior in the model, leading to diverse coding styles.
-One could hypothesize several reasons for these observed patterns.
-The propensity for conciseness at lower temperatures may reflect LLM’s training data, where probably concise solutions were more prevalent or deemed more ‘correct’.
-On the flip side, human developers, often juggling multiple considerations like future code extensions, comments for peer developers, or even personal coding style, might craft lengthier solutions.
-Furthermore, the narrowing of disparities at higher temperatures can be attributed to the model’s increased willingness to explore varied coding styles.
-At higher temperatures, the LLM possibly mimics a broader spectrum of human coding patterns, capturing the essence of diverse coding habits and styles found in its training corpus.
-Finding 1: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
+3.4.1 Token Frequency
+Table II: Top 50 tokens from human- and machine-authored code from CodeLlama
+Rank Human-Authored Tokens Machine-Authored Tokens
+1–10 . ( ) = ’ , : self ” [ . - , ( ) self : ” ’ if
+11–20 ] if return in for not None 0 1 == = return not [ def raise isinstance ] == path
+21–30 else + is name { } path data raise - name 0 __class__ __name__ None os { / } %
+31–40 try * os len format get and True value isinstance else TypeError str ‘ __init__ is ¿ // the
+41–50 args key % np i x kwargs except False or in 1 ; value kwargs #include + __str__ for ValueError
+Table II lists the top 50 tokens from human- and machine-authored code when $T$=0.2.
+Due to space limit, we omit the results when $T$=1.0, which has a similar result.
+From the results, we have several noteworthy observations:
+Common Tokens: Human- and machine-authored code shares a commonality in their usage of certain tokens, including punctuation marks such as “.”, “(”, “)”, and structural keywords such as “if”, “return”, and “else”.
+This is because LLMs acquire foundational coding syntax after being trained on extensive human-written code corpora.
+Error Handling: Tokens associated with error handling like “raise” and “TypeError”, are more prevalent in machine-authored code.
+This difference implies that machine programmers emphasize robustness and exception handling more explicitly.
+Programming Paradigms: Tokens indicating object-oriented programming like “self” and “__init__” are prominent in both human- and machine-authored code, which illustrates the model’s training alignment with this paradigm.
+However, machine-authored code appears to favor more boilerplate tokens like “__class__” and also “__name__”, which could stem from its training on diverse object-oriented codebase.
+Finding 1: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
+3.4.2 Syntax Element Distribution
+Figure 1: Syntax element distribution of the studied code corpus
+The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code.
+The “keyword”, “operator”, “syntactic symbols”, and “whitespace” proportions remain largely consistent between human and machine-authored code, across both temperatures.
+This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
+Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
+Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machines favor reusing established identifiers over creating new ones, possibly a reflection of its programming paradigm influenced by its training data.
+Literal: Machine-authored code consistently shows a marginally higher inclination towards literals.
+Across both temperatures, the machine exhibits an increase in the literal proportion as opposed to the human-written code.
+This indicates a potential proclivity of machine-authored code to process raw data more frequently.
+This could be attributed to the machine’s broader exposure to varied tasks involving data manipulations.
+Comment: Machine-authored code has much more comments when $T$=1.0.
+This observation hints machine’s increased emphasis on code documentation and explanation with higher temperatures, when it becomes less deterministic and more exploratory.
+Finding 2: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
+3.4.3 Zipf’s and Heaps’ Laws
 (a) Zipf’s Law ($T=0.2$)
@@
 (d) Heaps’ Law ($T=1.0$)
-Figure 2. Comparison of lexical diversity on machine- and human-authored code
-3.4.2. Lexical Diversity
+Figure 2: Comparison of Zipf’s and Heaps’ laws on machine- and human-authored code
 Figure 2 offers a comprehensive understanding of coding tendencies of both human and machine programmers.
@@
 Additionally, certain patterns within the training data might have been overemphasized due to model overfitting, leading the machine to a skewed preference.
-Finding 2: Machines demonstrate a preference for a limited spectrum of frequently-used tokens, whereas human code exhibits a richer diversity in token selection.
-3.4.3. Token Frequency
-Table 2. Top 50 tokens from human- and machine-authored code from CodeLlama on CodeSearchNet with $T=0.2$
-Rank Human-Authored Tokens Machine-Authored Tokens
-1–10 . ( ) = ’ , : self ” [ . - , ( ) self : ” ’ if
-11–20 ] if return in for not None 0 1 == = return not [ def raise isinstance ] == path
-21–30 else + is name { } path data raise - name 0 __class__ __name__ None os { / } %
-31–40 try * os len format get and True value isinstance else TypeError str ‘ __init__ is ¿ // the
-41–50 args key % np i x kwargs except False or in 1 ; value kwargs #include + __str__ for ValueError
-Table 2 lists the top 50 tokens from human- and machine-authored code when $T$=0.2.
-Due to space limit, we omit the results when $T=1.0$, which has a similar result.
-From the results, we have several noteworthy observations:
-Common Tokens: Human- and machine-authored code shares a commonality in their usage of certain tokens, including punctuation marks such as ‘.’, ‘(’, ‘)’, and structural keywords such as ‘if’, ‘return’, and ‘else’.
-This is because LLMs acquire foundational coding syntax after being trained on extensive human-written code corpora.
-Error Handling: Tokens associated with error handling like ‘raise’ and ‘TypeError’, are more prevalent in machine-authored code.
-This difference implies that machine programmers emphasize robustness and exception handling more explicitly.
-Programming Paradigms: Tokens indicating object-oriented programming like ‘self’ and ‘__init__’ are prominent in both human- and machine-authored code, which illustrates the model’s training alignment with this paradigm.
-However, machine-authored code appears to favor more boilerplate tokens like ‘__class__’ and also ‘__name__’, which could stem from its training on diverse object-oriented codebase.
-Finding 3: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
-3.4.4. Token Category Proportion
-Figure 3. Proportions of token categories in the studied code corpus
-The analysis of token category proportions, as visualized in Figure 3, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code.
-The ‘keyword’, ‘operator’, ‘syntactic symbols’, and ‘whitespace’ proportions remain largely consistent between human and machine-authored code, across both temperatures.
-This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
-Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers:
-Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machines favor reusing established identifiers over creating new ones, possibly a reflection of its programming paradigm influenced by its training data.
-Literal: Machine-authored code consistently shows a marginally higher inclination towards literals.
-Across both temperatures, the machine exhibits an increase in the literal proportion as opposed to the human-written code.
-This indicates a potential proclivity of machine-authored code to process raw data more frequently.
-This could be attributed to the machine’s broader exposure to varied tasks involving data manipulations.
-Comment: Machine-authored code has much more comments when $T$=1.0.
-This observation hints machine’s increased emphasis on code documentation and explanation with higher temperatures, when it becomes less deterministic and more exploratory.
-Finding 4: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
-3.4.5. Naturalness
-Figure 4. Distribution of naturalness scores for human- and machine-authored code
+Finding 3: Machines demonstrate a preference for a limited spectrum of frequently-used tokens, whereas human code exhibits a richer diversity in token selection.
+3.4.4 Number of Tokens and Lines
+(a) Number of tokens ($T=0.2$)
+(b) Number of tokens ($T=1.0$)
+(c) Number of lines ($T=0.2$)
+(d) Number of lines ($T=1.0$)
+Figure 3: Distribution of code length for machine- and human-authored code
+Figure 3 presents the distribution of code length under different settings.
+For the temperature setting of $0.2$, machine-authored code exhibits more conciseness, both in token and line numbers.
+As the temperature increases to $T=1.0$, we witness a convergence of distributions.
+The gap narrows, yet the machine’s preference for relatively concise code persists.
+This reveals that higher temperatures induce more exploratory generative behavior in the model, leading to diverse coding styles.
+One could hypothesize several reasons for these observed patterns.
+The propensity for conciseness at lower temperatures may reflect LLM’s training data, where probably concise solutions were more prevalent or deemed more “correct”.
+On the flip side, human developers, often juggling multiple considerations like future code extensions, comments for peer developers, or even personal coding style, might craft lengthier solutions.
+Furthermore, the narrowing of disparities at higher temperatures can be attributed to the model’s increased willingness to explore varied coding styles.
+At higher temperatures, the LLM possibly mimics a broader spectrum of human coding patterns, capturing the essence of diverse coding habits and styles found in its training corpus.
+Finding 4: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
+3.4.5 Token Likelihood and Rank
+Figure 4: Distribution of naturalness scores for machine- and human-authored code
 Figure 4 shows that there is a great discrepancy of naturalness between machine- and human-authored code.
 Compared to human-authored code, the log likelihood scores of machine-authored code are mostly higher and the log rank scores are mostly lower, indicating that machine’s code is more “natural” than human-written code.
-Such observation in source code is consistent with the findings in natural language (Gehrmann et al., 2019; Ippolito et al., 2020; Mitchell et al., 2023; Solaiman et al., 2019).
-Table 3. The naturalness of different token categories
+Such observation in source code is consistent with the findings in natural language [39, 38, 11, 40].
+Table III: The naturalness of different categories of syntax elements. Statistical significance $p<0.001$.
 Category Log Likelihood Log Rank
@@
 ALL -1.658 -0.827 0.831 0.319 0.811 0.492
-Table 3 summarises the comparison results in terms of each token category at $T=0.2$.
+Table III summarises the comparison results in terms of each token category at $T=0.2$.
 An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens.
-This highlights a distinctive aspect of coding practices: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns.
+This highlights a distinctive aspect of coding styles: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns.
 Humans, however, influenced by individual styles and practices, exhibit a wider variety in their use of whitespaces.
@@
 Finding 5: Machine-authored code exhibits higher “naturalness” than human-authored code, and the disparity is more pronounced in tokens such as comments, and whitespaces, which are more reflective of individual coding styles.
-4. Detecting Machine Generated Code
-The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces.
-This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those structural tokens that best characterize the machine’s preference.
+4 Detecting Machine Generated Code
+The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespaces.
+This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those stylistic tokens that best characterize the machine’s preference.
 Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-generated code.
-Our approach adapts the perturbation-based detection framework DetectGPT (Mitchell et al., 2023) to the code domain.
-4.1. Problem Formulation
+Our approach adapts the perturbation-based detection framework DetectGPT [11] to the code domain.
+4.1 Problem Formulation
 We formulate machine-generated code detection as a classification task that predicts whether a given code snippet $x$ is produced by a source model $p_{\theta}$.
@@
 We introduce the naturalness score and the perturbation strategy $q(\cdot|x)$ in our approach in the following sections.
-4.2. Measuring Naturalness
-Previous methods usually use the log likelihood of tokens to measure the naturalness of machine-authored content (Hindle et al., 2016; Solaiman et al., 2019).
-However, the log rank of tokens shows better performance comparing the naturalness of machine- and human-authored text (Mitchell et al., 2023; Gehrmann et al., 2019), because it offers a smoother and robust representation of token preference.
-Unlike DetectGPT which directly calculates the log likelihood of tokens, we adopt the Normalized Perturbed Log Rank (NPR) score (Su et al., 2023) to capture the naturalness.
-The NPR score is formally defined as:
-(3) $\mathbf{NPR}\left(x,p_{\theta},q\right)\triangleq\frac{\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log r_{\theta}\left(\tilde{x}\right)}{\log r_{\theta}(x)},$
+4.2 Measuring Naturalness
+Previous methods usually use the log likelihood of tokens to measure the naturalness of machine-authored content [32, 40].
+However, the log rank of tokens shows better performance comparing the naturalness of machine- and human-authored text [11, 39], because it offers a smoother and robust representation of token preference.
+Unlike DetectGPT which directly calculates the log likelihood of tokens, we adopt the Normalized Perturbed Log Rank (NPR) score [45] to capture the naturalness. The NPR score is formally defined as:
+$\mathbf{NPR}\left(x,p_{\theta},q\right)\triangleq\frac{\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log r_{\theta}\left(\tilde{x}\right)}{\log r_{\theta}(x)},$ (3)
 where $\log r_{\theta}(x)$ is the logarithm of the rank order of text $x$ sorted by likelihood under model $p_{\theta}$.
-In practice, $\mathbf{NPR}\left(x,p_{\theta},q\right)$ has been demonstrated to be more accurate for differentiating text origins, outperforming the log likelihood discrepancy $\mathbf{d}\left(x,p_{\theta},q\right)$ (Su et al., 2023).
-4.3. Perturbation Strategy
-Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in structural coding style.
-Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below.
-Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
-Space Insertion
-Let $C$ represent the set of all possible locations to insert spaces in a code segment.
-We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations.
-For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$.
-Mathematically, this can be represented as:
-(4) $n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$
-Newline Insertion
-We split the codes into lines and obtain a set $L$ of lines.
-A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations.
-For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$.
-Formally, this is expressed as:
-(5) $n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$
-We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations.
-Through this step, we instill randomness at a granular structural level, thereby amplifying the perturbation’s efficacy.
-Our perturbation strategy introduces several distinct advantages over the conventional methods (Mitchell et al., 2023; Su et al., 2023) using MLM to perturb the code, which will be discussed in Section 6.2.
+In practice, $\mathbf{NPR}\left(x,p_{\theta},q\right)$ has been demonstrated to be more accurate for differentiating text origins, outperforming the log likelihood discrepancy $\mathbf{d}\left(x,p_{\theta},q\right)$ [45].
 Data: code $x$, source model $\mathcal{M}$, number of perturbations $k$, decision threshold $\epsilon$, parameters $\alpha$, $\beta$, $\lambda_{\text{spaces}}$, and $\lambda_{\text{newlines}}$
@@
 3       if $p\leq 0.5$ then
-             // Spaces Insertion for Perturbed Sample
+             // Spaces Insertion
 4             Let $C$ represent the set of all possible locations to insert spaces in $x$;
@@
 12      else
-             // Newlines Insertion for Perturbed Sample
+             // Newlines Insertion
 13             Split the perturbed code $x$ into a set $L$ of lines;
@@
 24 end for
-25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\mathbf{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
-26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\mathbf{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
-27 Calculate variance: $\tilde{\sigma}_{x}^{2}\leftarrow\frac{1}{k-1}\sum_{i}\left(\mathbf{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)-\tilde{\mu}\right)^{2}$;
-28 if $\frac{\text{NPR}_{x}}{\sqrt{\tilde{\sigma}_{x}}}>\epsilon$ then
+25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
+26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
+[…]
        return true ;
-        // Probably machine-generated
-29
-30else
+        // Probably machine-authored
+28
+29else
        return false ;
-        // Probably human-written
-31
+        // Probably human-authored
+30
 31 end if
-Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Structured Code Perturbation
+Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Stylized Code Perturbation
+4.3 Perturbation Strategy
+Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in coding style.
+Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below.
+Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
+4.3.1 Space Insertion
+Let $C$ represent the set of all possible locations to insert spaces in a code segment.
+We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations.
+For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$.
+Mathematically, this can be represented as:
+$n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$ (4)
+4.3.2 Newline Insertion
+We split the codes into lines and obtain a set $L$ of lines.
+A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations.
+For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$.
+Formally, this is expressed as:
+$n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (5)
+We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations.
+Through this step, we instill randomness at a granular stylistic level, thereby amplifying the perturbation’s efficacy.
+Our perturbation strategy introduces several distinct advantages over the conventional methods [11, 45] using MLM to perturb the code, which will be discussed in Section 6.2.
 Algorithm 1 summarizes the entire workflow of DetectCodeGPT.
-Our algorithm harnesses structured code perturbation to differentiate between human- and machine-authored code.
-At the core of our approach is the strategic insertion of spaces (Lines 5-9) and newlines (Lines 11-16) in code, a process that simulates the inherent randomness in human coding habits.
-The algorithm operates by generating perturbed versions of the code and then evaluating their NPR scores (Lines 20-22) with respect to the source model $\mathcal{M}$.
-The threshold parameter $\epsilon$ in Line 23, pivotal for making the detection decision, offers flexibility in catering to different application scenarios.
+Our algorithm harnesses stylized code perturbation to differentiate between human- and machine-authored code.
+At the core of our approach is the strategic insertion of spaces (Lines 4-9) and newlines (Lines 11-16) in code, a process that simulates the inherent randomness in human coding styles.
+The algorithm operates by generating perturbed versions of the code and then evaluating their NPR scores (Lines 20-21) with respect to the source model $\mathcal{M}$.
+The threshold parameter $\epsilon$ in Line 22, pivotal for making the detection decision, offers flexibility in catering to different application scenarios.
 By adjusting $\epsilon$, users can balance between false positives and false negatives, tailoring the detection sensitivity according to the specific needs of the deployment context.
-5. Evaluation
+5 Evaluation
 We conduct experiments to evaluate the effectiveness of DetectCodeGPT, aiming to answer the following research questions.
@@
   • RQ4: How resilient is our method to cross-model code source detection?
-5.1. Datasets
-Apart from CodeSearchNet (Husain et al., 2020) described in Section 3.3, we select Python code from The Stack (Kocetkov et al., 2022) as another evaluation dataset.
+5.1 Datasets
+Apart from CodeSearchNet [21] described in Section 3.3, we select Python code from The Stack [22] as another evaluation dataset.
 Similar to CodeSearchNet, The Stack provides codes from a variety of open-source projects representative of real-world scenarios.
-We use a parsed and filtered version (Manh et al., 2023) of this dataset and also concatenate the function definitions with their corresponding comments as prompts.
+We use a parsed and filtered version [46] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1].
 For each combination of dataset and model, we construct 500 human and machine code pairs for evaluation.
 The maximum length of both human- and machine-authored code is trimmed to 128 tokens.
-5.2. Evaluation Metric
-Following prior works (Mitchell et al., 2023; Su et al., 2023), our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code.
+5.2 Evaluation Metric
+Following prior works [11, 45], our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code.
 Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as:
@@
 This makes the metric both interpretable and insightful, offering a clearer picture of the model’s discriminating capabilities.
-5.3. Baselines and Models
+5.3 Baselines and Models
 Our evaluation is benchmarked against a diverse range of zero-shot machine-generated text detection techniques, all of which leverage the predicted token-wise conditional distributions of the source model for detection.
 A supervised baseline is also included to demonstrate the effectiveness of our method.
-  • Log $p(x)$ (Solaiman et al., 2019): Utilizes the source model’s average token-wise log probability to gauge code naturalness. Machine-generated code tends to have a higher score.
-  • Entropy (Gehrmann et al., 2019): Interprets high average entropy in the model’s predictive distribution as indicative of machine generation.
-  • (Log-) Rank (Gehrmann et al., 2019; Ippolito et al., 2020): The average observed rank or log rank of each token in the LLM prediction, with machine-generated passages typically showing smaller average values.
-  • DetectGPT (Mitchell et al., 2023): Leverages the log probability of the original code and its perturbed variants to compute the perturbation discrepancy gap.
-  • DetectLLM (Su et al., 2023): Introduces two methods, one blends log likelihood with log rank to compute LRR, and the other improves DetectGPT by incorporating the NPR score.
-  • Supervised Method: a simple supervised detection method that predicts the authorship of a given code snippet using a binary classifier.
-Following OpenAI’s RoBERTa-based (Liu et al., 2019) GPT-2 detector [^2] https://github.com/openai/gpt-2-output-dataset/tree/master/detector, we employ CodeBERT (Feng et al., 2020) with an MLP head as the classifier and train it on a combination of 1000 samples generated by each model at each setting.
+  • Log $p(x)$ [40]: Utilizes the source model’s average token-wise log probability to gauge code naturalness. Machine-generated code tends to have a higher score.
+  • Entropy [39]: Interprets high average entropy in the model’s predictive distribution as indicative of machine generation.
+  • (Log-) Rank [39, 38]: The average observed rank or log rank of each token in the LLM prediction, with machine-generated passages typically showing smaller average values.
+  • DetectGPT [11]: Leverages the log probability of the original code and its perturbed variants to compute the perturbation discrepancy gap.
+  • DetectLLM [45]: Introduces two methods, one blends log likelihood with log rank to compute LRR, and the other improves DetectGPT by incorporating the NPR score.
+  • GPTSniffer [47]: a supervised baseline that trains CodeBERT [48] to predict the authorship of a given code snippet.
+Following OpenAI’s RoBERTa-based [49] GPT detector [^2] https://github.com/openai/gpt-2-output-dataset/tree/master/detector, we train the model on a combination of 1000 samples generated by each model at each setting.
 As for models to generate machine’s code, in consideration of GPU memory limitations, we judiciously selected a range of advanced LLMs with varying parameter counts, from 1 billion to 7 billion.
-We selected Incoder (Fried et al., 2022), Phi-1 (Gunasekar et al., 2023), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023), CodeGen2 (Nijkamp et al., 2023) and CodeLlama (Rozière et al., 2023) from Huggingface  [^3] https://huggingface.co/models.
+We selected Incoder [24], Phi-1 [50], StarCoder [51], WizardCoder [25], CodeGen2 [4] and CodeLlama [6] from Huggingface  [^3] https://huggingface.co/models.
 This diverse array of models, each trained on distinct datasets and boasting different capacities, offers a comprehensive spectrum for evaluating our method.
-5.4. Experimental Setup
-For code generation with different models, we adopted the top-$p$ sampling strategy (Holtzman et al., 2020) with $p$=$0.95$ following (Chen et al., 2021).
+5.4 Experimental Setup
+For code generation with different models, we adopted the top-$p$ sampling strategy [41] with $p$=$0.95$ following [1].
 We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section 3.2.
 The maximum length constraint for generated codes was set at $128$ tokens.
-With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model (Allal et al., 2023), revealed optimal values.
+With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values.
 Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}=3$ and $\lambda_{\text{newlines}}=2$.
 For all experiments, we maintained a consistent configuration of generating $50$ perturbations.
-For the DetectGPT and DetectLLM tasks, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model (Wang et al., 2023).
+For the DetectGPT and DetectLLM, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model [52].
+And as for the supervised baseline GPTSniffer, we trained the CodeBERT model for 5 epochs with a batch size of 16 and a learning rate of 2e-5 with AdamW optimizer.
 All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
-Table 4. Performance (AUROC) of various detection methods
+Table IV: Performance (AUROC) of various detection methods. Statistical significance $p<0.001$.
 Dataset Code LLM Detection Methods
-$\log p(x)$ Entropy Rank Log Rank  DetectGPT LRR NPR  Supervised  DetectCodeGPT
+$\log p(x)$ Entropy Rank Log Rank  DetectGPT LRR NPR  GPTSniffer  DetectCodeGPT
 CodeSearchNet ($T=0.2$) Incoder (1.3B) 0.9810 0.1102 0.8701 0.9892 0.4735 0.9693 0.8143 0.9426 0.9896
@@
 CodeLlama (7B) 0.5929 0.5260 0.6451 0.6091 0.6116 0.6365 0.6226 0.7494 0.6595
-5.5. Detection Performance (RQ2)
-Table 4 delineates the results of various methods.
+Average - 0.7649 0.3961 0.7228 0.7724 0.6357 0.7050 0.7178 0.6507 0.8308
+5.5 Detection Performance (RQ1)
+Table IV delineates the results of various methods.
 According to the results, DetectCodeGPT consistently outperforms baseline methods.
+Compared to the strongest baseline Log Rank, our method achieves an average relative improvement of 7.6% in AUROC.
 In an impressive 21 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama.
-The high AUROC values achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
+We also repeated the experiments 10 times and employed a paired t-test to assess the statistical significance of the performance differences between the methods, setting the significance level at 0.05.
+Results show that the performance superiority of our method was statistically significant, with p-values less than 0.001.
+The high AUROC scores achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
 We can observe that the challenge of detection notably increases at a temperature setting of $T$=$1.0$ than $T$=$0.2$.
-This is possibly due to the higher randomness at this temperature, where models are likely to generate outputs with greater diversity.
+This is possibly due to the higher randomness at this temperature, where models are likely to generate outputs with greater diversity in styles.
 Despite these increased difficulties, the proposed method maintains its leading position in detection accuracy.
-It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised approach.
+It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised GPTSniffer.
 The unstable performance of the supervised model reveals that it’s hard to detect machine-generated codes relying on the training data.
 The ability of our zero-shot approach to adapt and retain high performance under different levels of generative temperature variations is indicative of its sophisticated detection capability, making it a valuable tool for maintaining the integrity of codebases against the infiltration of machine-generated code in practical applications.
-5.6. Ablation Study (RQ3)
-Table 5. Performance of different perturbation strategies
+(a) Example 1
+(b) Example 2
+(c) Example 3
+(d) Example 4
+Figure 5: Examples of machine- and human-authored code snippets with corresponding predictions.
+To further illustrate the effectiveness of our approach, we present some representative examples of machine-generated code detection by DetectCodeGPT and two most competitive baselines in Figure 5.
+We set the mean score of all the code snippets as the decision threshold $\epsilon$ for each method.
+Examples 1 is a machine-generated code snippet.
+DetectCodeGPT can correctly detect it, while the baselines fail.
+In this example, the first and last “if” statements are separated from the rest of the code with newlines, and the two “if” statements in the middle are modularized together since they have similar functionalities.
+Such modularization and separation of code blocks are captured by DetectCodeGPT thanks to the stylized perturbation.
+Example 2 and 3 are human-written codes that are misclassified by other baselines, but DetectCodeGPT correctly identifies them.
+In Example 2, we can observe that the code blocks are sometimes separated with newlines (e.g., lines 5-7), but sometimes not (as seen with the rest of the code).
+In Example 3, although the code blocks are well separated with newlines, the human author omitted the spaces between operators “//” and “/”, but add spaces between “*” and “+”.
+Such freely and randomly stylized code reveals the inherent randomness in human coding habits.
+It’s hard for the baselines that only rely on token-wise conditional distributions to capture such randomness in coding styles.
+In contrast, DetectCodeGPT effectively leverages the style information to distinguish between machine-generated and human-written code again.
+These examples demonstrate the effectiveness of our method in detecting machine-generated code via stylized code perturbation.
+However, there are also some cases where DetectCodeGPT fails to detect machine-generated code.
+Example 4 is a human-written code misclassified as machine-generated by all the approaches.
+We can observe that this code snippet is well-structured with newlines and spaces, and the comments are well-aligned with the code.
+Its resemblance to machine-generated code is striking, posing a significant challenge for distinction.
+This example highlights the difficulty of detecting machine-generated code among well-structured human-written codes with standard coding styles.
+5.6 Ablation Study (RQ2)
+Table V: Performance of different perturbation strategies
 Perturb. Type MLM   Newline Space   Newline&Space
@@
 In our ablation study, we compared the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset.
-The results summarized in Table 5 illuminates the comparative advantage of our structured perturbation approach.
-We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM (Mitchell et al., 2023; Su et al., 2023) for natural language.
+The results summarized in Table V illuminates the comparative advantage of our stylistic perturbation approach.
+We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM [11, 45] for natural language.
 Also, the combination of newline and space perturbations further enhances the detection performance, with the highest AUROC score of 0.8852 at $T=0.2$ and 0.6660 at $T=1.0$.
 The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
-5.7. Impact of Perturbation Count (RQ4)
+5.7 Impact of Perturbation Count (RQ3)
 To gauge the impact of perturbation count on the efficacy of our method, we conducted experiments with varying numbers of perturbations.
-Table 6. Impact of varying the number of perturbations
+Table VI: Impact of varying the number of perturbations
 #Perturbations 10 20 50 100 200
@@
 $T=1.0$ 0.5558 0.6584 0.6660 0.6660 0.6662
-Results in Table 6 reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach.
+Results in Table VI reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach.
 Notably, an increase to 20 perturbations already yields robust detection performance, with further increments leading to diminishing improvements.
@@
 This implies that our method is not only effective but also resource-efficient.
-5.8. Performance of Cross-Model Detection (RQ5)
-Figure 5. Cross-model detection performance
+5.8 Performance of Cross-Model Detection (RQ4)
+Figure 6: Cross-model detection performance
 In previous sections, we primarily assessed the efficacy of DetectCodeGPT within a white-box framework, where the detection model can be the same as the source model.
@@
 And we can also notice a performance decrease when detecting CodeGen2’s output.
-This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text (Nijkamp et al., 2023).
-However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in (Mireshghallah et al., 2023).
-These results indicate that DetectCodeGPT is a model-free method that is robust against model discrepancies.
-6. Discussion
-In this section, we will analyze why our method is effective and discuss its strengths and limitations.
-6.1. Why is DetectCodeGPT Effective?
+This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text [4].
+However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in [53].
+These results indicate that DetectCodeGPT is a model-free method that is robust against model discrepancies, making it a viable solution for real-world applications where the source model could be unknown or inaccessible.
+6 Discussion
+6.1 Why is DetectCodeGPT Effective?
 We attribute the effectiveness of our DetectCodeGPT to the following two factors:
-Preservation of Code Correctness
+6.1.1 Preservation of Code Correctness
 While the perturbation method in DetectGPT masks continuous spans of text and requires another LLM to recover, such recovery often leads to structurally compromised code.
@@
 So our perturbation strategy ensures that the underlying structure and correctness of the code remain intact, thereby ensuring the effectiveness of our method.
-Emulation of Human Randomness
-As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of spaces and newlines than machines during coding.
-For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to structure the code in a more standardized and compact manner.
+6.1.2 Emulation of Human Randomness
+As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of stylistic tokens such as spaces and newlines than machines.
+For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to stylize the code in a more standardized and modularized manner.
 Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section 3.
-6.2. Strength of DetectCodeGPT
-DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding preferences.
+6.2 Strength of DetectCodeGPT
+DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding styles.
 We summarize the strengths of DetectCodeGPT as follows:
-Enhanced Efficiency
-Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection (Mitchell et al., 2023; Su et al., 2023), as verified in Section 5.7, perturbing code using MLM will lead to huge computational cost.
+6.2.1 Enhanced Efficiency
+Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection [11, 45], as verified in Section 5.7, perturbing code using masked language modeling with other pretrained models will lead to huge computational cost.
 On the contrary, our method will save much more time and bring more efficiency.
@@
 Mathematically, if $T_{\text{DetectGPT}}$ and $T_{\text{Ours}}$ denote the computational times of DetectGPT and our method respectively, we know that $T_{\text{Ours}}\ll T_{\text{DetectGPT}}$.
-Superior Zero-Shot Performance
+6.2.2 Superior Zero-Shot Performance
 DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets.
-This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated in Section 5.8.
+This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated and verified in Section 5.8.
 In comparison to the supervised methods, which may suffer from data drift or the need for constant updates, our approach is not only more adaptable but also delivers better performance across various data sources.
 This makes DetectCodeGPT an agile and reliable tool for machine-generated code detection in a variety of contexts.
-6.3. Limitations and Future Directions
+6.3 Limitations and Future Directions
 The main limitations of our work lie in the following two aspects: Firstly, due to the time and computational constraints, we only focus on a set of LLMs within 7B parameters for both empirical analysis and the machine-generated code detection task.
@@
 Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness, such as $T$=$1.0$.
-We note from Table 4 that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement.
-Although ensembling multiple detection models may help to improve the detection performance (Mireshghallah et al., 2023), we look forward to explore more effective perturbation strategies based on code structure to further enhance the detection efficacy.
-Additionally, the development of watermarking strategies (Sun et al., 2023) for machine-generated code represents an intriguing avenue for future research.
+We note from Table IV that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement.
+Although ensembling multiple detection models may help to improve the detection performance [53], we look forward to exploring more effective perturbation strategies based on code style to further enhance the detection efficacy.
+Additionally, the development of watermarking strategies [54] for machine-generated code represents an intriguing avenue for future research.
 Given that our method can effectively identify machine’s code, it could be adapted to assist in the creation of watermarking techniques.
 For instance, embedding a watermark could be achieved by inserting newlines and spaces in a code-specific pattern, which would increase detectablility without compromising functionality.
-Moreover, our work also highlights the need for the development of more robust detection methods against possible attacks (Hu et al., 2023; Krishna et al., 2023).
-As discussed in Section 3, an important difference between machine- and human-authored code is naturalness.
-We can possibly leverage the insertion of whitespace to imitate a less ‘natural’ code style and thereby make machine’s code more difficult to detect.
-This potential strategy highlights the need for the development of more robust detection methods against possible attacks (Hu et al., 2023; Krishna et al., 2023).
-7. Related Work
-Recently, there has been much effort in detecting machine-generated text (Yang et al., 2023; Wu et al., 2023b).
-The two main categories of detection methods are zero-shot and training-based methods, and our DetectCodeGPT falls into the former category.
-As for zero-shot methods, they are usually based on the discrepancy between likelihood and rank information of human and machine’s texts (Gehrmann et al., 2019; Ippolito et al., 2020; Solaiman et al., 2019).
-Leveraging the hypothesis in DetectGPT (Mitchell et al., 2023) that machine-generated text often has a negative curvature in the log probability when the text is perturbed, many perturbation based methods have been proposed (Mitchell et al., 2023; Su et al., 2023; Bao et al., 2023).
-These methods usually perturb the text by masking a continuous span of tokens and then recover the perturbed text using another LLM like T5 (Ahmad et al., 2021).
+7 Related Work
+Recently, there has been much effort in detecting machine-generated text [9, 55].
+The two main categories of detection methods are zero-shot and training-based methods, and our DetectCodeGPT falls into the former category, which eliminates the need for training data and brings more generalization ability.
+As for zero-shot methods, they are usually based on the discrepancy between likelihood and rank information of human and machine’s texts [39, 38, 40].
+Leveraging the hypothesis in DetectGPT [11] that machine-generated text often has a negative curvature in the log probability when the text is perturbed, many perturbation based methods have been proposed [11, 45, 56].
+These methods usually perturb the text by masking a continuous span of tokens and then recover the perturbed text using another LLM like T5 [29].
 The benefit of these methods is that they are zero-shot and can be applied to any LLM without access to training data.
-However, the perturbation process is time-consuming and computationally expensive,
-When it comes to the training-based methods, fine-tuning the RoBERTa(Liu et al., 2019) or T5 (Ahmad et al., 2021) model with data collected from different model families at different decoding settings is a common practice (Tian et al., 2023; Zhan et al., 2023; Chen et al., 2023).
-Additional information like graph structure (Liu et al., 2022), perplexity from proxy models (Wu et al., 2023a) have been shown to be helpful for detection.
-Moreover, techniques like adversarial training (Hu et al., 2023) and contrastive learning (Liu et al., 2022) have also been proposed to improve the detection performance.
-The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model.
-However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text (Lee et al., 2023).
-The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation (Lee et al., 2023; Sun et al., 2023).
+However, the perturbation process is time-consuming and computationally expensive.
+When it comes to the training-based methods, fine-tuning the RoBERTa[49] or T5 [29] model with data collected from different model families at different decoding settings is a common practice [15, 57, 58].
+Additional information like graph structure [59], perplexity from proxy models [60] have been shown to be helpful for detection.
+Moreover, techniques like adversarial training [61] and contrastive learning [59] have also been proposed to improve the detection performance.
+The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model [11].
+However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text [13].
+GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [47].
+And the predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation [13, 54].
 The detection of these watermarks subsequently enables the recognition of code generated by machines.
-It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism (Cox et al., 2002; Collberg and Thomborson, 1999).
+It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism [62, 63].
 Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
-8. Conclusion
-In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the age of AI.
-By pioneering a comparative analysis across five different attributes, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces.
-Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the structural nuances of code with a novel perturbation strategy that is simple yet effective.
+8 Conclusion
+In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the era of AI.
+By pioneering a comparative analysis across three aspects of code, namely, naturalness, style, and correctness, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylized tokens such as whitespaces.
+Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the stylized nuances of code with a novel perturbation strategy that is simple yet effective.
 The empirical success of DetectCodeGPT against several benchmarks confirms its potential to bolster security measures in programming and establish a more symbiotic relationship between human developers and LLMs.

### v3

@@
-Index Terms: Machine-generated code detection, large language models, code generation, empirical analysis
+Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers
 Abstract
@@
 Thus, its applicability falters when applied to code.
-In this paper, we carefully study the specific patterns that characterize machine and human-authored code.
+In this paper, we carefully study the specific patterns that characterize machine- and human-authored code.
 Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source.
 We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance.
-Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct stylized patterns of code.
+Based on our findings, we propose DetectCodeGPT, a novel method for detecting machine-generated code, which improves DetectGPT by capturing the distinct stylized patterns of code.
 Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency.
-Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
-Index Terms:
-Machine-generated code detection, large language models, code generation, empirical analysis
-1 Introduction
-The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect.
-LLMs trained on code [3, 4, 5, 6], sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation [7, 8].
-Their ability to generate syntactically correct and functionally robust code challenges traditional dynamics and introduces a new era of efficiency and innovation in software creation, maintenance, and evolution.
-While capable of generating human-like programs, LLMs bring ambiguity of whether a software artifact is created by a human or machine, causing integrity and authenticity issues in software development.
-This indistinction can lead to various challenges, such as the misattribution of code ownership for bug/fault triage and potential vulnerabilities in machine-generated code that may go unnoticed due to overreliance on its perceived robustness.
-Furthermore, the ease of generating code could lead to inflated assessments of project workloads, skewing economic valuations and potentially compromising the reliability of software.
-This blending of human and machine efforts not only raises questions about the trustworthiness of the software but also threatens the integrity of coding as a discipline, wherein the true authorship and the effort invested in creating software artifacts become obscured.
+Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code  [^1] Our code and data are available at https://github.com/YerbaPage/DetectCodeGPT.
+I Introduction
+The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering tasks such as code generation.
+Through extensive training on ultra-large code corpora [3, 4, 5, 6], LLMs acquire the ability to generate syntactically and functionally correct code, bringing a new era of efficiency and innovation in software creation, maintenance, and evolution.
+While capable of generating human-like code, LLMs bring ambiguity of whether a software artifact is created by human or machine, causing integrity and authenticity issues in software development.
+This indistinction can lead to various challenges, such as misattribution of code ownership for bug triage and inflated assessments of developer workloads.
+Potential vulnerabilities in machine-generated code may go unnoticed due to the overreliance on its perceived robustness.
+The blending of human and machine efforts not only raises questions about the trustworthiness of the software but also threatens the integrity of development process, wherein the true authorship and the effort invested in creating software artifacts become obscured.
 Addressing these concerns is pivotal in maintaining a transparent and secure software development lifecycle.
 Recently, there has been a growing research trend in detecting machine-generated texts [7, 8].
-Perturbation-based methods like DetectGPT [11], have achieved state-of-the-art results in identifying machine-generated text.
-DetectGPT employs likelihood score discrepancies between the original text and its perturbed variants to enhance detection accuracy.
-And the perturbation process is performed by a pretrained language model (e.g., T5 [12]) to mask portions of the original text and then reconstruct it.
-While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid syntactic constraints.
-Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code [13].
-In this paper, we conduct a comparative analysis of the patterns between machine- and human-authored code from three aspects, including lexical diversity, conciseness, and naturalness.
-Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespace tokens.
-Capitalizing on our findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code.
-Our DetectCodeGPT improves DetectGPT by capturing the distinct styles of code.
-Diverging from conventional techniques that depend on external pretrained model (e.g.
-CodeT5 [14]) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code by strategically inserting spaces and newlines.
-Eschewing the need for an LLM for perturbations, our method provides much higher efficiency.
-Since the proposed perturbation strategy well captures the stylistic distinctions between machine- and human-authored code, our DetectCodeGPT also ensures higher efficacy.
-To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models, ranging from 1.3 to 7 billion parameters, and operating at two different decoding conditions.
-The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward [9, 15], detecting machine-authored code presents a unique challenge.
-Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement.
+Perturbation-based methods like DetectGPT [9] have achieved state-of-the-art results in identifying machine-generated text.
+These methods employ likelihood score discrepancies between the original text and its various LLM-perturbed variants for detection, capturing the distinct patterns of machine-generated text that machines tend to prefer to a smaller set of phrases and expressions.
+However, such detection methods tailored for natural language texts face challenges when applied to code, as code requires strict adherence to syntactic rules, while natural language can maintain coherence with more variation [10].
+This situation highlights a significant gap in existing research: a lack of in-depth assessment of the intrinsic features of machine- and human-authored code, crucial for understanding the unique patterns of machine-generated code and devising effective detection methods.
+In this paper, we conduct a comparative analysis of the distinct patterns between machine- and human-authored code from three aspects, including lexical diversity, conciseness, and naturalness.
+Through our analysis, we uncover that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens and adhere to common programming paradigms, and the disparity is more pronounced in stylistic tokens such as the whitespace tokens.
+Based on the findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code.
+We extend the perturbation-based framework of DetectGPT by strategically inserting stylistic tokens, such as whitespace and newline characters, to capture the distinct patterns between machine- and human-authored code.
+This approach capitalizes on our observation that the disparity in coding styles is more pronounced in these stylistic tokens.
+By directly manipulating the code, DetectCodeGPT eliminates the need for an external pre-trained model, thereby enhancing both efficiency and effectiveness in the detection process.
+To evaluate the effectiveness of DetectCodeGPT, we have conducted extensive experiments on two datasets across six code language models.
+The results demonstrate that DetectCodeGPT significantly outperforms the state-of-the-art methods by 7.6% in terms of AUC.
+Moreover, it proves to be a model-free and robust method against model discrepancies, making it viable for real-world applications with unknown or inaccessible source models.
 Our contributions can be summarized as follows:
-  • To our knowledge, we are the first to conduct a comprehensive and thorough analysis of the differences between machine- and human-authored code.
-Our study sheds light on essential insights that further advance the utility of LLMs in coding.
-  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct stylistic patterns of code.
+  • To our knowledge, we are the first to conduct a comprehensive and thorough analysis of the distinct patterns of LLM-generated code.
+Our study sheds light on essential insights that can further advance the utility of LLMs in programming.
+  • We propose a novel method for detecting machine-generated code by leveraging its distinct stylistic patterns.
   • We extensively evaluate the DetectCodeGPT across a variety of settings and show the effectiveness of our approach.
-2 Background
-2.1 Large Language Models for Code
-GPT language models [16, 17] based on Transformer [18] decoder has achieved remarkable success in natural language processing tasks [19].
-In the file of code generation, Codex [1] and AlphaCode [20] are pioneering works to train large language models on code.
-These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens [16].
-Mathematically, this is captured as:
-$\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$ (1)
-where $x_{i}$ denotes each instance in the training dataset and $p_{\theta}(x_{i})$ represents the model’s probability distribution parameterized by $\theta$.
-And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub [21, 22, 23].
+II Background
+II-A Large Language Models for Code
+Large language models [11, 12] based on Transformer [13] decoder has achieved remarkable success in natural language processing tasks [14].
+In the domain of code generation, Codex [1] and AlphaCode [15] are pioneering works to train large language models on code.
+The training data often contain millions of code in different programming languages collected from open source repositories like GitHub [16, 17, 18].
 Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle [19, 20, 4, 5] and also instruction fine-tuning [21, 6].
-And recent large language models pretrained on a mixture of code and natural language like ChatGPT [26], GPT4 [27] and LLaMA [28] has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
-2.2 Perturbation-Based Detection of Machine-Generated Text
+Recent large language models pretrained on a mixture of programming and natural languages like ChatGPT [22] and LLaMA [23] have also shown promising results on code generation tasks.
+II-B Perturbation-Based Detection of Machine-Generated Text
 In the realm of machine-generated text detection, perturbation based method like DetectGPT [9] stands as the state-of-the-art technology [7].
@@
 This method effectively captures the more nuanced and variable nature of human-generated text compared to the more formulaic and patterned output of language models.
-3 Empirical Analysis
+III Empirical Analysis
 In this section, we conduct a comparative analysis of the distinct features of machine- and human-authored code.
-3.1 Study Design
+III-A Study Design
 To gain insights into distinctions between human and machine programmers, we consider three primary aspects that are relevant to coding styles, namely diversity, conciseness, and naturalness [25, 26, 27, 28, 29], which can be measured by specific metrics.
-3.1.1 Lexical Diversity
+III-A1 Lexical Diversity
 Lexical diversity indicates the richness and variety of vocabulary present in a corpus.
@@
 The attribute indicates the core vocabulary utilized by human and machine programmers, shedding light on their coding preferences and tendencies.
-Syntax Element Distribution refers to the proportion of syntax elements (e.g., keywords, identifiers) in the code corpus.
-Understanding the distribution of syntax elements in code is akin to dissecting the anatomy of a language.
-It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human- and machine-authored code.
-To delve into the syntax element distribution, we analyze code with tree-sitter [^1] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table I.
-We then compute the proportion of each category in the code corpus.
-Zipf’s and Heaps’ Laws were initially identified in natural languages [35, 36], and later verified in the scope of programming languages [30, 31].
-Zipf’s law states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ [35].
-In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed.
-Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
-We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human lexical usage.
 Table I: Studied categories of Python code tokens
@@
 […]
-3.1.2 Conciseness
+Syntax Element Distribution refers to the proportion of syntax elements (e.g., keywords, identifiers) in the code corpus.
+Understanding the distribution of syntax elements in code is akin to dissecting the anatomy of a language.
+It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human- and machine-authored code.
+To delve into the syntax element distribution, we analyze code with tree-sitter [^2] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table I.
+We then compute the proportion of each category in the code corpus.
+Zipf’s and Heaps’ Laws were initially identified in natural languages [30, 31], and later verified in the scope of programming languages [25, 26].
+Zipf’s law states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ [30].
+In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed.
+Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
+We investigate how closely machine-authored code aligns with Zipf’s and Heaps’ laws compared to human-authored code, which could reflect the models’ ability to mimic human’s lexical usage.
+III-A2 Conciseness
 Conciseness stands as a cornerstone attribute when characterizing code [32, 28, 29].
@@
 Number of lines helps us understand organizational choices, as spreading code across more lines can reflect a focus on readability and structure [29].
-3.1.3 Naturalness
+III-A3 Naturalness
 The concept of code naturalness suggests that programming languages share a similar degree of regularity and predictability with natural languages [27].
@@
 To pinpoint the code elements that most significantly affect the score discrepancies, we also present the mean scores on different syntax element categories in Table I for comparison.
-3.2 Experimental Setup
+Table II: Top 50 tokens from human- and machine-authored code from CodeLlama
+Rank Human-Authored Tokens Machine-Authored Tokens
+1–10 . ( ) = ’ , : self ” [ . - , ( ) self : ” ’ if
+11–20 ] if return in for not None 0 1 == = return not [ def raise isinstance ] == path
+21–30 else + is name { } path data raise - name 0 __class__ __name__ None os { / } %
+[…]
+41–50 args key % np i x kwargs except False or in 1 ; value kwargs #include + __str__ for ValueError
+III-B Experimental Setup
 We choose the state-of-the-art CodeLlama model [6] to generate code.
@@
 All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
-3.3 Dataset Preparation
-To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus [21], a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase.
-We used function signatures and their accompanying comments as prompts for the model as in [1], and we collected the corresponding bodies of these functions to represent human-written code.
+III-C Dataset Preparation
+To compare with human-authored code, we extract 10,000 Python functions randomly from the CodeSearchNet corpus [16], which is curated from a wide range of open-source GitHub projects.
+We use the function signatures and their accompanying comments as prompts for the model as in [1].
+We also collect the corresponding bodies of these functions to represent human-written code.
 While acknowledging that current models, including CodeLlama and even ChatGPT [38], may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet [39, 38], the choice of this dataset is deliberate and insightful.
-Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval [1] or MBPP [44] dataset, offers a more representative assessment.
-It allows us to analyze the differences between human and machine’s code when faced with broader, practical applications.
-3.4 Results and Analysis
-We present the results and analysis regarding each code attribute introduced in Section 3.1.
-3.4.1 Token Frequency
-Table II: Top 50 tokens from human- and machine-authored code from CodeLlama
-Rank Human-Authored Tokens Machine-Authored Tokens
-1–10 . ( ) = ’ , : self ” [ . - , ( ) self : ” ’ if
-11–20 ] if return in for not None 0 1 == = return not [ def raise isinstance ] == path
-21–30 else + is name { } path data raise - name 0 __class__ __name__ None os { / } %
-31–40 try * os len format get and True value isinstance else TypeError str ‘ __init__ is ¿ // the
-41–50 args key % np i x kwargs except False or in 1 ; value kwargs #include + __str__ for ValueError
+Challenging the models against various real-world project code rather than simple programming problems, akin to those in the HumanEval [1] or MBPP [40] dataset, offers a more representative assessment.
+It allows us to analyze the differences between human- and machine-authored code when faced with broader, practical applications.
+III-D Results and Analysis
+We present the results and analysis regarding each code attribute introduced in Section III-A.
+III-D1 Token Frequency
 Table II lists the top 50 tokens from human- and machine-authored code when $T$=0.2.
@@
 However, machine-authored code appears to favor more boilerplate tokens like “__class__” and also “__name__”, which could stem from its training on diverse object-oriented codebase.
-Finding 1: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
-3.4.2 Syntax Element Distribution
-Figure 1: Syntax element distribution of the studied code corpus
+Finding 1: Machine-authored code pays more attention to exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming paradigms.
+III-D2 Syntax Element Distribution
+Figure 1: Syntax element distribution of the code corpus
 The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code.
@@
 Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
-Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machines favor reusing established identifiers over creating new ones, possibly a reflection of its programming paradigm influenced by its training data.
-Literal: Machine-authored code consistently shows a marginally higher inclination towards literals.
-Across both temperatures, the machine exhibits an increase in the literal proportion as opposed to the human-written code.
-This indicates a potential proclivity of machine-authored code to process raw data more frequently.
-This could be attributed to the machine’s broader exposure to varied tasks involving data manipulations.
+Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machine-authored code may have a more compact style with fewer identifiers.
+Literal: Machine-authored code consistently shows a slightly higher tendency towards using literals.
+Across both temperatures, the machine code exhibits an increase in the literal proportion compared to the human-written code.
+This suggests that machine-authored code may be more likely to process raw data directly.
+This could result from the machine’s training on diverse data manipulation tasks.
 Comment: Machine-authored code has much more comments when $T$=1.0.
 This observation hints machine’s increased emphasis on code documentation and explanation with higher temperatures, when it becomes less deterministic and more exploratory.
-Finding 2: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
-3.4.3 Zipf’s and Heaps’ Laws
+Finding 2: Machine-authored code tends to use fewer identifiers, more literals for direct data processing, and have more comments when the generation temperature grows.
+III-D3 Zipf’s and Heaps’ Laws
 (a) Zipf’s Law ($T=0.2$)
@@
 Finding 3: Machines demonstrate a preference for a limited spectrum of frequently-used tokens, whereas human code exhibits a richer diversity in token selection.
-3.4.4 Number of Tokens and Lines
+III-D4 Number of Tokens and Lines
 (a) Number of tokens ($T=0.2$)
@@
 At higher temperatures, the LLM possibly mimics a broader spectrum of human coding patterns, capturing the essence of diverse coding habits and styles found in its training corpus.
-Finding 4: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
-3.4.5 Token Likelihood and Rank
-Figure 4: Distribution of naturalness scores for machine- and human-authored code
+Finding 4: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their stylistic preferences.
+III-D5 Token Likelihood and Rank
+Figure 4: Distribution of naturalness scores
 Figure 4 shows that there is a great discrepancy of naturalness between machine- and human-authored code.
@@
 Finding 5: Machine-authored code exhibits higher “naturalness” than human-authored code, and the disparity is more pronounced in tokens such as comments, and whitespaces, which are more reflective of individual coding styles.
-4 Detecting Machine Generated Code
-The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespaces.
-This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those stylistic tokens that best characterize the machine’s preference.
-Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-generated code.
-Our approach adapts the perturbation-based detection framework DetectGPT [11] to the code domain.
-4.1 Problem Formulation
-We formulate machine-generated code detection as a classification task that predicts whether a given code snippet $x$ is produced by a source model $p_{\theta}$.
+IV Detecting Machine Generated Code
+The empirical results suggest that machines tend to write more concise and natural code with a narrower spectrum of tokens, adhering to common programming paradigms, and the disparity is more pronounced in stylistic tokens such as whitespaces.
+This sparks a new idea for detecting machine-authored code: instead of perturbing arbitrary tokens, we focus on perturbing those stylistic tokens that best characterize the machine’s preference.
+Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-authored code.
+IV-A Problem Formulation
+We formulate the detection of machine-authored code as a classification task, which predict whether or not a given code snippet $x$ is produced by a source model $p_{\theta}$.
 For this purpose, we transform $x$ to an equivalent form $\tilde{x}$ through a perturbation process $q(\cdot|x)$.
@@
 We introduce the naturalness score and the perturbation strategy $q(\cdot|x)$ in our approach in the following sections.
-4.2 Measuring Naturalness
+IV-B Measuring Naturalness
 Previous methods usually use the log likelihood of tokens to measure the naturalness of machine-authored content [27, 35].
@@
 In practice, $\mathbf{NPR}\left(x,p_{\theta},q\right)$ has been demonstrated to be more accurate for differentiating text origins, outperforming the log likelihood discrepancy $\mathbf{d}\left(x,p_{\theta},q\right)$ [41].
+IV-C Perturbation Strategy
+Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in coding style.
+Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below.
+Detailed explanations on the effectiveness of these perturbations are given in Section VI-A.
+IV-C1 Space Insertion
+Let $C$ represent the set of all possible locations to insert spaces in a code segment.
+We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations.
+For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$.
+Mathematically, this can be represented as:
+$n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$ (3)
+IV-C2 Newline Insertion
+We split the code into lines and obtain a set $L$ of lines.
+A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations.
+For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$.
+Formally, this is expressed as:
+$n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (4)
 Data: code $x$, source model $\mathcal{M}$, number of perturbations $k$, decision threshold $\epsilon$, parameters $\alpha$, $\beta$, $\lambda_{\text{spaces}}$, and $\lambda_{\text{newlines}}$
-// Detection using NPR score on Perturbed Samples
 1 for $i\leftarrow 1$ to $k$ do
@@
              // Spaces Insertion
-4             Let $C$ represent the set of all possible locations to insert spaces in $x$;
+4             Let $C$ represent all possible locations to insert spaces in $x$;
 5             Select $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$;
@@
 24 end for
-25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
-26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
+25Estimate NPR: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
 […]
@@
         // Probably machine-authored
-28
-29else
+27
+28else
        return false ;
         // Probably human-authored
-30
+29
 30 end if
 Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Stylized Code Perturbation
-4.3 Perturbation Strategy
-Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in coding style.
-Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below.
-Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
-4.3.1 Space Insertion
-Let $C$ represent the set of all possible locations to insert spaces in a code segment.
-We randomly select a subset $C_{s}\subseteq C$ such that $|C_{s}|=\alpha\times|C|$, where $\alpha\in[0,1]$ is a fraction representing the code locations.
-For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$.
-Mathematically, this can be represented as:
-$n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$ (4)
-4.3.2 Newline Insertion
-We split the codes into lines and obtain a set $L$ of lines.
-A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations.
-For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$.
-Formally, this is expressed as:
-$n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (5)
 We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations.
 Through this step, we instill randomness at a granular stylistic level, thereby amplifying the perturbation’s efficacy.
-Our perturbation strategy introduces several distinct advantages over the conventional methods [11, 45] using MLM to perturb the code, which will be discussed in Section 6.2.
+Our perturbation strategy introduces several distinct advantages over the conventional methods [9, 41] using MLM to perturb the code, which will be discussed in Section VI-B.
 Algorithm 1 summarizes the entire workflow of DetectCodeGPT.
@@
 By adjusting $\epsilon$, users can balance between false positives and false negatives, tailoring the detection sensitivity according to the specific needs of the deployment context.
-5 Evaluation
+V Evaluation
 We conduct experiments to evaluate the effectiveness of DetectCodeGPT, aiming to answer the following research questions.
@@
   • RQ3: What is the impact of varying the number of perturbations on the detection performance?
-  • RQ4: How resilient is our method to cross-model code source detection?
-5.1 Datasets
-Apart from CodeSearchNet [21] described in Section 3.3, we select Python code from The Stack [22] as another evaluation dataset.
-Similar to CodeSearchNet, The Stack provides codes from a variety of open-source projects representative of real-world scenarios.
+  • RQ4: How effective is our method in cross-model code detection?
+V-A Datasets
+Apart from CodeSearchNet [16] described in Section III-C, we select Python code from The Stack [17] as another evaluation dataset.
+Similar to CodeSearchNet, The Stack provides code from a variety of open-source projects representative of real-world scenarios.
 We use a parsed and filtered version [42] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1].
-For each combination of dataset and model, we construct 500 human and machine code pairs for evaluation.
-The maximum length of both human- and machine-authored code is trimmed to 128 tokens.
-5.2 Evaluation Metric
-Following prior works [11, 45], our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code.
+For each combination of dataset and model, we sample 500 human and machine code pairs for evaluation.
+The maximum length of code is trimmed to 128 tokens.
+V-B Studied Models
+We investigate machine-generated code by a diverse array of advanced LLMs, including Incoder [20], Phi-1 [43], StarCoder [44], WizardCoder [21], CodeGen2 [4] and CodeLlama [6].
+We obtain their checkpoints from Huggingface  [^3] https://huggingface.co/models with different parameter sizes (1B-7B).
+V-C Evaluation Metric
+Following prior works [9, 41], our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC).
 Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as:
@@
 This makes the metric both interpretable and insightful, offering a clearer picture of the model’s discriminating capabilities.
-5.3 Baselines and Models
-Our evaluation is benchmarked against a diverse range of zero-shot machine-generated text detection techniques, all of which leverage the predicted token-wise conditional distributions of the source model for detection.
-A supervised baseline is also included to demonstrate the effectiveness of our method.
+V-D Baselines
+Our evaluation is benchmarked against a diverse range of zero-shot machine-generated text detection techniques.
+And a supervised baseline is also included to demonstrate the advantages of our zero-shot method:
   • Log $p(x)$ [35]: Utilizes the source model’s average token-wise log probability to gauge code naturalness. Machine-generated code tends to have a higher score.
@@
 Following OpenAI’s RoBERTa-based [47] GPT detector [^4] https://github.com/openai/gpt-2-output-dataset/tree/master/detector, we train the model on a combination of 1000 samples generated by each model at each setting.
-As for models to generate machine’s code, in consideration of GPU memory limitations, we judiciously selected a range of advanced LLMs with varying parameter counts, from 1 billion to 7 billion.
-We selected Incoder [24], Phi-1 [50], StarCoder [51], WizardCoder [25], CodeGen2 [4] and CodeLlama [6] from Huggingface  [^3] https://huggingface.co/models.
-This diverse array of models, each trained on distinct datasets and boasting different capacities, offers a comprehensive spectrum for evaluating our method.
-5.4 Experimental Setup
+V-E Experimental Setup
 For code generation with different models, we adopted the top-$p$ sampling strategy [36] with $p$=$0.95$ following [1].
-We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section 3.2.
-The maximum length constraint for generated codes was set at $128$ tokens.
+We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section III-B.
+The maximum length constraint for generated code was set at $128$ tokens.
 With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values.
@@
 Average - 0.7649 0.3961 0.7228 0.7724 0.6357 0.7050 0.7178 0.6507 0.8308
-5.5 Detection Performance (RQ1)
+V-F Detection Performance (RQ1)
 Table IV delineates the results of various methods.
@@
 Despite these increased difficulties, the proposed method maintains its leading position in detection accuracy.
-It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised GPTSniffer.
-The unstable performance of the supervised model reveals that it’s hard to detect machine-generated codes relying on the training data.
-The ability of our zero-shot approach to adapt and retain high performance under different levels of generative temperature variations is indicative of its sophisticated detection capability, making it a valuable tool for maintaining the integrity of codebases against the infiltration of machine-generated code in practical applications.
+It is worth mentioning that our zero-shot framework often outperforms the supervised GPTSniffer, highlighting the challenges of detecting machine-generated code with training data-dependent supervised models.
+The stable performance across various generation settings showcases our method’s advanced detection capabilities.
+This makes it an effective solution in practical applications.
+To further illustrate the effectiveness of our approach, we present some representative examples by DetectCodeGPT and two most competitive baselines in Figure 5, using decision thresholds $\epsilon$ set at the mean scores of all code snippets.
+Examples 1 is a machine-generated code snippet.
+DetectCodeGPT can correctly detect it, while the baselines fail.
+In this example, the first and last “if” statements are separated from the rest of the code with newlines, and the two “if” statements in the middle are modularized together since they have similar functionalities.
+Such modularization and separation of code blocks are captured by DetectCodeGPT thanks to the stylized perturbation.
+Example 2 and 3 are human-written code that are misclassified by other baselines, but DetectCodeGPT correctly identifies them.
+In Example 2, we can observe that the code blocks are sometimes separated with newlines (e.g., lines 5-7), but sometimes not (as seen with the rest of the code).
+In Example 3, although the code blocks are well separated with newlines, the human author omitted the spaces between operators “//” and “/”, but add spaces between “*” and “+”.
+Such freely and randomly stylized code reveals the inherent randomness in human coding habits.
+Baselines, relying solely on token-wise conditional distributions, struggle to capture the coding style’s randomness, whereas DetectCodeGPT effectively utilizes style information to discern between machine-generated and human-written code, demonstrating its prowess in detection through stylized code perturbation.
+However, there are also cases where DetectCodeGPT fails.
+Example 4 is a human-written code misclassified as machine-generated by all the approaches.
+We can observe that this code snippet is well-structured with newlines and spaces.
+Its resemblance to machine-generated code is striking, posing a significant challenge for distinction.
+This example highlights the difficulty of detecting machine-generated code among well-structured human-written code with standard coding styles.
 (a) Example 1
@@
 Figure 5: Examples of machine- and human-authored code snippets with corresponding predictions.
-To further illustrate the effectiveness of our approach, we present some representative examples of machine-generated code detection by DetectCodeGPT and two most competitive baselines in Figure 5.
-We set the mean score of all the code snippets as the decision threshold $\epsilon$ for each method.
-Examples 1 is a machine-generated code snippet.
-DetectCodeGPT can correctly detect it, while the baselines fail.
-In this example, the first and last “if” statements are separated from the rest of the code with newlines, and the two “if” statements in the middle are modularized together since they have similar functionalities.
-Such modularization and separation of code blocks are captured by DetectCodeGPT thanks to the stylized perturbation.
-Example 2 and 3 are human-written codes that are misclassified by other baselines, but DetectCodeGPT correctly identifies them.
-In Example 2, we can observe that the code blocks are sometimes separated with newlines (e.g., lines 5-7), but sometimes not (as seen with the rest of the code).
-In Example 3, although the code blocks are well separated with newlines, the human author omitted the spaces between operators “//” and “/”, but add spaces between “*” and “+”.
-Such freely and randomly stylized code reveals the inherent randomness in human coding habits.
-It’s hard for the baselines that only rely on token-wise conditional distributions to capture such randomness in coding styles.
-In contrast, DetectCodeGPT effectively leverages the style information to distinguish between machine-generated and human-written code again.
-These examples demonstrate the effectiveness of our method in detecting machine-generated code via stylized code perturbation.
-However, there are also some cases where DetectCodeGPT fails to detect machine-generated code.
-Example 4 is a human-written code misclassified as machine-generated by all the approaches.
-We can observe that this code snippet is well-structured with newlines and spaces, and the comments are well-aligned with the code.
-Its resemblance to machine-generated code is striking, posing a significant challenge for distinction.
-This example highlights the difficulty of detecting machine-generated code among well-structured human-written codes with standard coding styles.
-5.6 Ablation Study (RQ2)
+V-G Ablation Study (RQ2)
 Table V: Performance of different perturbation strategies
@@
 $T=1.0$ 0.6226 0.6453 0.6504 0.6660
-In our ablation study, we compared the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset.
+In our ablation study, we compare the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset.
 The results summarized in Table V illuminates the comparative advantage of our stylistic perturbation approach.
@@
 The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
-5.7 Impact of Perturbation Count (RQ3)
-To gauge the impact of perturbation count on the efficacy of our method, we conducted experiments with varying numbers of perturbations.
+V-H Impact of Perturbation Count (RQ3)
+To gauge the impact of perturbation count on the efficacy of our method, we conduct experiments with varying numbers of perturbations.
 Table VI: Impact of varying the number of perturbations
@@
 This suggests that our method requires a relatively small number of perturbations to effectively discern between human- and machine-authored code.
-This implies that our method is not only effective but also resource-efficient.
-5.8 Performance of Cross-Model Detection (RQ4)
+This implies that our method is not only effective but also efficient.
+V-I Performance of Cross-Model Code Detection (RQ4)
 Figure 6: Cross-model detection performance
-In previous sections, we primarily assessed the efficacy of DetectCodeGPT within a white-box framework, where the detection model can be the same as the source model.
-Nonetheless, in real-world applications, it is often infeasible to access to the original code generation model for detection.
-To simulate such conditions, we engaged in cross-model detection experiments, wherein distinct models were utilized to compute the naturalness scores for detecting code generated by others.
-These evaluations were carried out using The Stack dataset at a temperature of 0.2.
-The results presented in Figure 6 highlight DetectCodeGPT’s adaptability in cross-model detection.
-While the algorithm excels in the white-box setting, its performance endures with only a slight reduction when subjected to cross-model application.
+In previous sections, we primarily assess the efficacy of DetectCodeGPT within a white-box framework, where we can get access to the logits of the original code generation model, as defined in Section IV-A.
+In real-world scenarios, accessing the original model for code generation detection is often impractical, so we conduct cross-model detection experiments, where we use other LLMs as surrogate models to compute the naturalness scores.
+The evaluation results on The Stack dataset at $T=0.2$ presented in Figure 6 highlight DetectCodeGPT’s adaptability in cross-model detection.
+While the algorithm excels in the white-box setting, its performance endures with only a slight reduction in cross-model application.
 For instance, StarCoder, when detecting code generated by WizardCoder and CodeLlama, yields AUROC scores of 0.92 and 0.87, respectively, compared to an AUROC of 0.93 when detecting its own output.
-And we can also notice a performance decrease when detecting CodeGen2’s output.
+We can also notice a performance decrease when detecting CodeGen2’s output.
 This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text [4].
@@
 These results indicate that DetectCodeGPT is a model-free method that is robust against model discrepancies, making it a viable solution for real-world applications where the source model could be unknown or inaccessible.
-6 Discussion
-6.1 Why is DetectCodeGPT Effective?
+VI Discussion
+VI-A Why is DetectCodeGPT Effective?
 We attribute the effectiveness of our DetectCodeGPT to the following two factors:
-6.1.1 Preservation of Code Correctness
-While the perturbation method in DetectGPT masks continuous spans of text and requires another LLM to recover, such recovery often leads to structurally compromised code.
-The unique syntax and constraints inherent to programming languages mean that even minor misinterpretations (e.g., misuse of an identifier) can render the code non-functional.
-The originally masked spans could be the only correct solution to recover the code and the recovered version from LLM bring minor mistakes easily.
-Such code-cracking perturbations will possibly lead to both a decrease in naturalness scores for human’s and machine’s codes, violating the assumption that the perturbation will have minimal impact on the code’s naturalness score if it is human-written in Section 2.2.
-In contrast, inserting newlines and spaces does not affect the correctness of the code in most cases.
-So our perturbation strategy ensures that the underlying structure and correctness of the code remain intact, thereby ensuring the effectiveness of our method.
-6.1.2 Emulation of Human Randomness
-As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of stylistic tokens such as spaces and newlines than machines.
+VI-A1 Preservation of Code Correctness
+The mask-and-recover perturbation in DetectGPT brings minor mistakes easily (e.g., misuse of an identifier), rendering the code non-functional and has negative impact on the naturalness score.
+Such code-cracking perturbations will violate the assumption of minimal impact on the code’s naturalness score if it is human-written in Section II-B.
+In contrast, inserting newlines and spaces does not affect the correctness of the code in most cases, thereby ensuring the effectiveness of our method.
+VI-A2 Emulation of Human Randomness
+As discussed in Section III-D5, human inherently exhibit less naturalness and more randomness in their use of stylistic tokens such as spaces and newlines than machines.
 For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to stylize the code in a more standardized and modularized manner.
-Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section 3.
-6.2 Strength of DetectCodeGPT
-DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding styles.
-We summarize the strengths of DetectCodeGPT as follows:
-6.2.1 Enhanced Efficiency
-Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection [11, 45], as verified in Section 5.7, perturbing code using masked language modeling with other pretrained models will lead to huge computational cost.
-On the contrary, our method will save much more time and bring more efficiency.
-This translates to a significant conservation of computational resources, thus expediting the detection process and enhancing scalability.
-Mathematically, if $T_{\text{DetectGPT}}$ and $T_{\text{Ours}}$ denote the computational times of DetectGPT and our method respectively, we know that $T_{\text{Ours}}\ll T_{\text{DetectGPT}}$.
-6.2.2 Superior Zero-Shot Performance
-DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets.
-This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated and verified in Section 5.8.
-In comparison to the supervised methods, which may suffer from data drift or the need for constant updates, our approach is not only more adaptable but also delivers better performance across various data sources.
-This makes DetectCodeGPT an agile and reliable tool for machine-generated code detection in a variety of contexts.
-6.3 Limitations and Future Directions
-The main limitations of our work lie in the following two aspects: Firstly, due to the time and computational constraints, we only focus on a set of LLMs within 7B parameters for both empirical analysis and the machine-generated code detection task.
+Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section III.
+VI-B Strength of DetectCodeGPT
+Compared with existing methods, DetectCodeGPT eliminates the need to perturb code multiple times for each LLM and thus brings more efficiency.
+Compared with supervised counterparts, DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets.
+This model-agnostic advantage means that it can be generalized across various code LLMs.
+VI-C Limitations and Future Directions
+The main limitations of our work lie in the following two aspects: Firstly, due to the computational constraints, we only focus on a set of LLMs within 7B parameters.
 As the landscape of LLMs rapidly evolves, incorporating a wider array of more and larger LLMs could significantly bolster the generalizability and robustness of our findings.
-Secondly, our current analysis centers exclusively on Python code, while the features of other programming languages may not be fully explored and the effectiveness of our method on other languages is not verified.
-However, based on the analysis of our method in Section 6.2, we believe that our method can be effectively generalized to other programming languages, especially where the functionality of code won’t be much affected after inserting newlines and spaces like C/C++, Java, and JavaScript.
-Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness, such as $T$=$1.0$.
+Secondly, our current analysis centers exclusively on Python code, while the features of other programming languages may not be fully explored.
+However, based on the analysis of our method in Section VI-B, we believe that our method can be effectively generalized to other languages, especially where the functionality of code won’t be much affected after inserting newlines and spaces like C/C++, Java, and JavaScript.
+Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness.
 We note from Table IV that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement.
-Although ensembling multiple detection models may help to improve the detection performance [53], we look forward to exploring more effective perturbation strategies based on code style to further enhance the detection efficacy.
-Additionally, the development of watermarking strategies [54] for machine-generated code represents an intriguing avenue for future research.
-Given that our method can effectively identify machine’s code, it could be adapted to assist in the creation of watermarking techniques.
-For instance, embedding a watermark could be achieved by inserting newlines and spaces in a code-specific pattern, which would increase detectablility without compromising functionality.
-7 Related Work
+Although ensembling multiple detection models may help improve the detection performance [49], we look forward to exploring more effective perturbation strategies based on code styles to further enhance the detection efficacy.
+VII Related Work
 Recently, there has been much effort in detecting machine-generated text [7, 50].
@@
 The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model [9].
-However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text [13].
+However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-generated text [10].
 GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [45].
-And the predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation [13, 54].
+The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training or generation [10, 58, 59].
 The detection of these watermarks subsequently enables the recognition of code generated by machines.
@@
 Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
-8 Conclusion
-In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the era of AI.
-By pioneering a comparative analysis across three aspects of code, namely, naturalness, style, and correctness, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylized tokens such as whitespaces.
-Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the stylized nuances of code with a novel perturbation strategy that is simple yet effective.
-The empirical success of DetectCodeGPT against several benchmarks confirms its potential to bolster security measures in programming and establish a more symbiotic relationship between human developers and LLMs.
-Our work marks a significant step towards responsible AI use in software engineering, ensuring clarity in code authorship and maintaining the integrity of code as a craft.
-Data Availability
-All the experimental data and source code used in this work are available at https://github.com/YerbaPage/DetectCodeGPT.
+VIII Conclusion
+In this paper, we perform an in-depth analysis of the nuanced differences between machine- and human-authored code across three aspects of code including lexical diversity, conciseness, and naturalness.
+The results provide new insights that machines tend to write more concise and natural code, adhering to common programming paradigms, and the disparity is more pronounced in stylized tokens such as whitespaces that represent the syntactic segmentation of code.
+Based on these insights, we have proposed a new detection method, DetectCodeGPT, which introduces a novel stylized perturbation strategy that is simple yet effective.
+The experimental results of DetectCodeGPT confirm its effectiveness, demonstrating its potential to help maintain the authorship and integrity of code.

### v5

@@
 Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency.
-Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code  [^1] Our code and data are available at https://github.com/YerbaPage/DetectCodeGPT.
+Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
+[^1] Code available at https://github.com/YerbaPage/DetectCodeGPT.
+Index Terms:
+machine-generated code detection, large language models, code generation, empirical analysis
 I Introduction
+footnotetext: * Xiaodong Gu is the corresponding author
 The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering tasks such as code generation.
@@
 Category Tree-sitter Types
-keyword
-def, return, else, if, for, while, $\ldots$
-identifier
-identifier, type_identifier
-literal
-string_content, integer, true, false, $\ldots$
-operator
-<, >, =, ==, +, $\ldots$
-syntactic symbol
-comment
-comment
-whitespace
+keyword def, return, else, if, for, while, $\ldots$
+identifier identifier, type_identifier
+literal string_content, integer, true, false, $\ldots$
+operator <, >, =, ==, +, $\ldots$
+syntactic symbol :, ), ], [, (, ,, ", , {, }, .
+comment comment
 […]
@@
 Since current LLMs on code are usually evaluated across different decoding temperatures [5, 37, 1, 6], we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings.
-The maximum length of the generated code is set to 512 tokens.
+The maximum length of the generated code is set to 512 tokens based on the memory constraints and the length distribution of human-written code.
 All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@
 This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
-Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
+Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significant from Chi-square test with $p<0.01$):
 Identifier: The identifiers constitute a significantly lower proportion among machine-authored code across both temperatures, indicating that machine-authored code may have a more compact style with fewer identifiers.
@@
 Figure 2 offers a comprehensive understanding of coding tendencies of both human and machine programmers.
-Starting with Zipf’s law, Figure 1(a) and 1(b) both delineate similar trends for human and machine programmers, corroborating the law’s applicability.
+Starting with Zipf’s law, Figure 2(a) and 2(b) both delineate similar trends for human and machine programmers, corroborating the law’s applicability.
 And we can observe machine’s heightened proclivity towards tokens ranked between 10 and 100, especially at $T$=$0.2$.
-Turning our attention to Heaps’ law, the near-linear trends in Figure 1(c)-1(d) reaffirm the law’s validity.
+Turning our attention to Heaps’ law, the near-linear trends in Figure 2(c)-2(d) reaffirm the law’s validity.
 Also, there’s a noticeable shallowness in the slope for machine’s code at $T$=$0.2$ revealing machine’s decreased lexical diversity.
@@
 whitespace -1.131 -2.740 1.609 0.429 1.441 1.012
-ALL -1.658 -0.827 0.831 0.319 0.811 0.492
-Table III summarises the comparison results in terms of each token category at $T=0.2$.
+ALL -0.827 -1.658 0.831 0.319 0.811 0.492
+Table VIII summarises the comparison results in terms of each token category at $T=0.2$.
 An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens.
@@
 For each location $c\in C_{s}$, we introduce a variable number of spaces, $n_{\text{spaces}}(c)$, which is drawn from a Poisson distribution $\mathcal{P}(\lambda_{\text{spaces}})$.
+The Poisson distribution is chosen to simulate the randomness in human coding styles, similar to the random text infilling strategy in BART [42].
 Mathematically, this can be represented as:
@@
 A subset $L_{n}\subseteq L$ is then chosen randomly, where $|L_{n}|=\beta\times|L|$, with $\beta\in[0,1]$ denoting the proportion of the line locations.
-For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$.
-Formally, this is expressed as:
+For each line $l\in L_{n}$, we introduce a variable number of newlines, $n_{\text{newlines}}(l)$, also sampled from a Poisson distribution $\mathcal{P}(\lambda_{\text{newlines}})$:
 $n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (4)
 Data: code $x$, source model $\mathcal{M}$, number of perturbations $k$, decision threshold $\epsilon$, parameters $\alpha$, $\beta$, $\lambda_{\text{spaces}}$, and $\lambda_{\text{newlines}}$
-1 for $i\leftarrow 1$ to $k$ do
+1
+2for $i\leftarrow 1$ to $k$ do
        // Random decision for type of perturbation
@@
 9                   Insert $n_{\text{spaces}}(c)$ spaces at location $c$ in $x$;
-9
+10
 11             end for
-11
+12
 13      else
@@
 18                   Insert $n_{\text{newlines}}(l)$ newlines after line $l$ in $x$;
-18
+19
 20             end for
-20
+21
 22       end if
 23      Store the perturbed code as $\tilde{x}_{i}$;
-23
+24
 25 end for
-25Estimate NPR: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
-[…]
+26
+27Estimate NPR: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
+28
+29if $\text{NPR}_{x}>\epsilon$ then
        return true ;
         // Probably machine-authored
-27
-28else
+30
+31else
        return false ;
         // Probably human-authored
-29
+32
 33 end if
+34
 Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Stylized Code Perturbation
@@
 V-A Datasets
-Apart from CodeSearchNet [16] described in Section III-C, we select Python code from The Stack [17] as another evaluation dataset.
+We carefully use a different split of the CodeSearchNet dataset [16] from the one used in the empirical study for evaluation.
+We select Python code from The Stack [17] as another evaluation dataset.
 Similar to CodeSearchNet, The Stack provides code from a variety of open-source projects representative of real-world scenarios.
@@
 WizardCoder (3B) 0.8293 0.3459 0.7484 0.8223 0.6377 0.6436 0.7929 0.7766 0.8384
-CodeGen2 (3.7B) 0.4816 0.6046 0.5631 0.4956 0.4337 0.5740 0.5178 0.4265 0.6660
-CodeLlama (7B) 0.5929 0.5260 0.6451 0.6091 0.6116 0.6365 0.6226 0.7494 0.6595
+CodeGen2 (3.7B) 0.4816 0.6046 0.5631 0.4956 0.4337 0.5740 0.5178 0.4265 0.6595
+CodeLlama (7B) 0.5929 0.5260 0.6451 0.6091 0.6116 0.6365 0.6226 0.7494 0.6660
 Average - 0.7649 0.3961 0.7228 0.7724 0.6357 0.7050 0.7178 0.6507 0.8308
@@
 Compared to the strongest baseline Log Rank, our method achieves an average relative improvement of 7.6% in AUROC.
-In an impressive 21 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama.
-We also repeated the experiments 10 times and employed a paired t-test to assess the statistical significance of the performance differences between the methods, setting the significance level at 0.05.
+In an impressive 20 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama.
+We also repeated the experiments 10 times and employed a Wilcoxon rank sum test to assess the statistical significance of the performance differences between the methods.
 Results show that the performance superiority of our method was statistically significant, with p-values less than 0.001.
@@
 VII Related Work
+VII-A Machine-Generated Text Detection
 Recently, there has been much effort in detecting machine-generated text [7, 51].
@@
 The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model [9].
-However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-generated text [10].
+VII-B Machine-Generated Code Detection
+Research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-generated text, according to the empirical study in [59].
 GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [46].
-The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training or generation [10, 58, 59].
+Concurrent works [60] and [61] also explore perturbation-based methods for detecting machine-generated code, under similar framework to DetectGPT for text [9].
+Our approach differs from these methods in that we perform a comprehensive empirical analysis of the differences between machine- and human-authored code.
+Based on the insights from the analysis, we proposed a innovative stylized perturbation strategy to achieve a more efficient and effective detection method.
+Another related topic revolve around code watermarking techniques, which embed unique markers into the code either during the training or generation [10, 62, 63].
 The detection of these watermarks subsequently enables the recognition of code generated by machines.
@@
 The experimental results of DetectCodeGPT confirm its effectiveness, demonstrating its potential to help maintain the authorship and integrity of code.
+Acknowledgement
+This research is supported by the National Key Research and Development Program of China (Grant No.
+2023YFB4503802) and the National Natural Science Foundation of China (Grant No.
+62102244).
+We would like to thank the anonymous reviewers for their valuable feedback and suggestions.
+Appendix
+Some results are omitted in the main text due to space constraints. We provide these additional results and analyses in this appendix.
+I Corresponding Results for Empirical Analysis
+The top tokens from machine-authored code at $T=1.0$ are shown in Table VII, corresponding to the results in Table II in the main text.
+Similar attention on exception handling and object-oriented programming tokens can also be observed at $T=1.0$.
+Table VII: Top tokens from machine-authored code at $T=1.0$
+Rank Machine-Authored Tokens
+1-10 . ’ ( ) self : , = ” if
+11-20 return def [ ] None in == not is import
+21-30 for { raise __name__ from } - data else get
+31-40 elif value path isinstance True + os ValueError key text
+41-50 x type 0 1 len append replace str @ f
+Figure 7 and Table VIII provide the distribution of naturalness scores and the naturalness of different categories of syntax elements, respectively.
+They correspond to the empirical analysis in Figure 4 and Table VIII in the main text.
+We can observe from Figure 7 that the trend of naturalness scores is similar to that at $T=0.2$.
+although the overlap between machine- and human-authored code is greater.
+The naturalness of different categories of syntax elements in Table VIII shows that whitespace tokens is still the most effective feature as in $T=0.2$.
+Figure 7: Distribution of naturalness scores with $T=1.0$
+Table VIII: The naturalness of different categories of syntax elements with $T=1.0$
+Category Log Likelihood Log Rank
+Machine Human $\Delta$ Machine Human $\Delta$
+keyword -2.025 -2.128 0.103 0.968 1.053 0.085
+identifier -0.787 -0.874 0.087 0.328 0.378 0.050
+literal -1.059 -1.364 0.305 0.454 0.630 0.176
+operator -1.614 -1.835 0.221 0.733 0.872 0.139
+symbol -0.968 -1.639 0.671 0.321 0.781 0.460
+comment -2.395 -3.028 0.633 1.180 1.610 0.430
+whitespace -2.058 -2.740 0.682 0.946 1.441 0.495
+ALL -1.367 -1.658 0.291 0.618 0.811 0.193
+II Cross-model Detection Performance at $T=1.0$
+Figure 8 illustrates the cross-model detection performance at $T=1.0$, corresponding to the results in Figure 6 in the main text.
+The results show that the cross-model detection performance of DetectCodeGPT has a similar trend across different temperatures.
+Figure 8: Cross-model detection performance at $T=1.0$
+III Impact of Varying Length for Detection
+The maximum length of code is trimmed to 128 tokens in our experiments.
+Figure 9 illustrates the trend in detection performance as the length of code snippets varies across different temperatures, compared against two of the most competitive baselines.
+This experiment uses the Stack dataset with CodeLlama.
+The results show that trimming the first 128 tokens of the snippets is sufficient to maintain high detection performance.
+(a) AUROC at $T=0.2$
+(b) AUROC at $T=1.0$
+Figure 9: AUROC with different code trimming length